### v4

# A Review on Knowledge Graphs for Healthcare: Resources, Applications, and Promises

Abstract.
Healthcare knowledge graphs (HKGs) are valuable tools for organizing biomedical concepts and their relationships. The recent advance of large language models (LLMs) has paved the way for building more comprehensive and accurate HKGs. This, in turn, can improve the reliability and evaluation of LLMs. However, the challenges and opportunities of HKGs are not fully understood, highlighting the need for detailed reviews. This work provides the first comprehensive review of HKGs, summarizing the pipeline and key techniques for HKG construction and successful HKG utilization in various health-related applications. Lastly, we highlight the opportunities for HKGs in the era of LLMs.
knowledge graph, healthcare, language models, multimodality, interpretability, trustworthy AI
$c$s: Applied computing Health informaticsccs: Applied computing Health care information systemsccs: Applied computing Bioinformaticsccs: Information systems Graph-based database modelsccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
Figure 1. The content overview of this review on healthcare knowledge graph.
A knowledge graph (KG) is a data structure that captures the relationships between different entities and their attributes (Ji et al., 2021; Nicholson and Greene, 2020). KG models and integrates data from various sources, including structured and unstructured data, and has been studied to support a wide range of applications such as search engines (Wang et al., 2019b), recommendation systems (Wang et al., 2019a; Zhou et al., 2020), and question answering (Lin et al., 2019; Yasunaga et al., 2021; Yan et al., 2021; Kan et al., 2021). Healthcare Knowledge Graph (HKG) facilitates an interpretable representation of medical concepts, e.g., drugs and disease, as well as the relations among those medical concepts. This data structure enables the connection of contexts and enhances clinical research and decision-making (Santos et al., 2022; Chandak et al., 2023).
On the data side, HKG is usually built on complex medical systems such as electronic health records, medical literature, clinical guidelines, and patient-generated data (Bouayad et al., 2017; Rajkomar et al., 2018). However, these data resources are often heterogeneous and distributed, which makes it challenging to integrate and analyze them effectively (Mehta and Pandit, 2018). This data heterogeneity can also lead to incomplete or inconsistent data representations, limiting their usefulness for downstream healthcare tasks (Dash et al., 2019). Additionally, the current use of domain-specific knowledge graphs may result in limited coverage and granularity of the knowledge captured across different levels. This hinders identifying correlations and relationships between medical concepts from multiple domains. These challenges highlight the need for continued research on HKGs to realize their full potential.
On the modeling side, the construction of HKGs can be done either from scratch or by integrating existing dataset resources. Many crucial steps, such as entity and relation extraction, can be optimized with natural language processing tools and algorithms. Recently, there have been significant advancements in general domain knowledge extraction, thanks to pre-trained large language models (LLMs) such as BERT (Devlin et al., 2019), GPT Series (Brown et al., 2020), and others. These models revolutionize the field and make it possible to integrate heterogeneous medical data from various sources effectively. The use of pre-trained models has also led to the development of more accurate and comprehensive medical ontologies and taxonomies (Zhang et al., 2021d; Yu et al., 2020; Wang et al., 2021; Zeng et al., 2021; Xu et al., 2020). This allows for the evaluation of generated contents from LLMs and reduces LLM hallucination.
A comprehensive healthcare knowledge graph has the potential to contribute to health research across various levels (Gyrard et al., 2018; Santos et al., 2022; Li et al., 2020). At the micro-scientific level, HKGs can help researchers identify new phenotypic and genotypic correlations and understand the underlying mechanisms of disease (Hassani-Pak and Rawlings, 2017), leading to more targeted and effective treatments (Seneviratne et al., 2021; Chandak et al., 2023). At the clinical care level, HKGs can be used to develop clinical decision support systems that provide clinicians with relevant information, improving clinical workflows and patient outcomes (Eberhardt et al., 2012; Castaneda et al., 2015). Therefore, a thorough review of existing literature on HKGs becomes an essential roadmap and invaluable resource to drive transformative advancements in the field.
This review is the first comprehensive overview of HKGs, covering contents shown in Figure 1. Specifically, in Section 2, we delve into the construction pipelines of HKGs, including both building from scratch and integration approaches, and highlight the key techniques employed in HKG construction. Additionally, we explore two standard utilization methods of HKGs, namely model-free and model-based approaches. In Section 3, we compile a comprehensive summary of existing HKG resources with their scopes and applications. Furthermore, Section 4 investigates the literature on mainstream health applications, offering an in-depth overview of the diverse use cases of HKGs in healthcare. Finally, we address promising research opportunities in the era of LLMs in Section 5.
2. Backgrounds
2.1. Knowledge Graphs for Healthcare (HKG)
Definition. A healthcare knowledge graph (HKG) is a domain-specific knowledge graph designed to capture medical concepts such as drugs, diseases, genes, phenotypes, and so on, and their relationships in a structured and semantic way.
Terminology. Our focus is on healthcare knowledge graphs (HKGs), which are structured, semantic representations of medical concepts and relationships. We also include ontologies and knowledge bases, which are commonly used in constructing HKGs. Ontologies define a set of concepts and categories in a domain, as well as the relationships between them, while knowledge bases store factual information about entities and their attributes. By covering these categories of terminology, we provide a comprehensive overview of the different types of resources available for organizing and representing medical knowledge in a structured and semantically rich manner.
[…]
L1/.style=draw=black, text width=3cm, L2/.style=draw=black, text width=4cm, L3/.style=draw=black, text width=5cm, L4/.style=draw=black, text width=5cm, {forest} for tree=draw,rounded corners,grow’=0,text width=2cm, text centered,edge+=-¿, forked edges, [Background, fill=orange!30, L1 [HKG Definition, fill=orange!20, L2] [HKG Construction, fill=orange!20, L2 [Construct from Scratch, fill=orange!10, L3] [Construct by Integration, fill=orange!10, L3] ] [HKG Utilization, fill=orange!20, L2 [Model-free Utilization, fill=orange!10, L3] [Model-based Utilization, fill=orange!10, L3] ] ]
Figure 2. Detailed taxonomy of the background section of healthcare knowledge graphs.
2.2. HKG Construction
Healthcare knowledge graphs can be constructed from scratch or through the integration of existing data resources.
Develop schemaCollect dataExtract knowledge triplesNormalize entities and relationsInfer missing linksUpdate and validate
Figure 3. The pipeline of constructing HKGs from scratch.
Constructing HKGs from Scratch. A multi-step pipeline, as in Figure 3, is used to construct HKGs from scratch.
  1. (1)
    The first step is to identify the scope and objectives. In most cases, researchers develop a schema (Guarino et al., 2009; Blagec et al., 2022) or use existing schemas (Guarino et al., 2009; Ashburner et al., 2000; Schriml et al., 2012; Bard et al., 2005) to serve as the formal and explicit specification of a domain, thus ensuring consistent, coherent, and aligned domain knowledge. Unlike the general domain KG, utilizing schemas is a common practice in HKG construction.
  2. (2)
    Secondly, researchers gather data from various sources, including medical literature, clinical trials, and patient-generated data. It is essential to ensure the quality and consistency of the data and to remove identifiable information for patient privacy.
  3. (3)
    The third step is to extract and transform the data into a structured format. This step involves identifying medical entities and creating relationships between them via specialized biomedical Natural Language Processing (NLP) tools (Song et al., 2021; Xing et al., 2020; Hahn and Oleynik, 2020).
  4. (4)
    Next, researchers map the entities and relationships to the chosen ontologies with the help of thesauruses (Bodenreider, 2004) or terminologies (Donnelly et al., 2006; Hirsch et al., 2016). This step ensures that the knowledge graph is interoperable with other healthcare systems and facilitates data integration.
  5. (5)
    Until now, an initial KG has been built. The next step is to populate the KG to infer missing links between entities. This inference can be done using graph databases (Wang et al., 2020) or link prediction models (Bordes et al., 2013; Lu and Yang, 2022).
  6. (6)
    The final step is to continuously update and validate the KG to ensure accuracy and relevance. This step involves incorporating new data and knowledge, refining the schema, and evaluating the quality of the KG.
Constructing HKGs by Integration. Considering significant efforts have been paid to construct and curate HKGs from scratch, it is promising to integrate these data resources to avoid repetitive work. Healthcare KG integration (also called Healthcare KG fusion) refers to the processing of merging two or more HKGs into a single, more comprehensive graph (Himmelstein et al., 2017; Su et al., 2023; Youn et al., 2022). The integration process is challenging because different HKGs may use different terminologies, schemas, or data formats. To address these challenges, researchers have developed various techniques and algorithms for knowledge graph fusion, including ontology matching (Faria et al., 2014; He et al., 2022), schema alignment (Suchanek et al., 2011; Maaroufi et al., 2014), entity resolution (Bachman et al., 2018; Hu et al., 2021), and conflicts resolution (Ma et al., 2023). These methods aim to identify and reconcile the differences between KGs.
Techniques for HKG Constructions. Traditionally, each step of HKG construction involves one specially designated model. For instance, Hidden Markov Models and Recurrent Neural Networks are widely used for healthcare named entity recognition, relation extraction, and other sequence tagging tasks, while Translational Models and Graph Neural Networks are used for HKG completion and conflicts resolution tasks. Recently, large language models (LLMs) have shown great utility to serve as a uniform tool for constructing KGs (Ye et al., 2022). Several key steps of constructing KGs, such as named entity recognition (Liang et al., 2020; Chen et al., 2023; Huang et al., 2022; Liu et al., 2022b), relation extraction (Zhuang et al., 2022; Lu et al., 2022; Yang et al., 2022), entity linking (De Cao et al., [n. d.]; Mrini et al., 2022; Cho et al., 2022), and KG completion (geng2022disentangled; Saxena et al., 2022; Xie et al., 2022; Shen et al., 2022), have been successfully tackled by these large foundation models. Early explorations of construction HKG with large foundation models show that healthcare entity normalization (Zhang et al., 2023; Agrawal et al., 2022), healthcare entity recognition (Fries et al., 2021; Hu et al., 2023), healthcare entity linking (Zhu et al., 2022b), and healthcare knowledge fusion (Lu et al., 2023) can also be performed, without extensive training on expensive healthcare annotated corpus. On the other hand, researchers start to construct KGs under the open-world assumption (Shi and Weninger, 2018; Das et al., 2020; Niu et al., 2021; Li et al., 2022; Lu and Yang, 2022), thus getting rid of the dependency on pre-defined schemas and exhaustive entity&relation normalization. Although open-world KGs greatly increase the coverage, ensuring the quality of extracted knowledge is still an open research challenge, especially for explainable and trustworthy HKGs.
Table 1. Resource of existing healthcare knowledge graph (HKG).
Name Node Types Edge Types Statistic Application
HetioNet (Himmelstein and Baranzini, 2015) 11 (e.g., drug, disease) 24 (e.g., drug-disease) #N: 47.0 K, #E: 2.3 M Medicinal Chemistry
DrKG (Ioannidis et al., 2020) 13 (e.g., disease, gene) 107 (e.g., disease-gene) #N: 97 K, #E: 5.8 M Medicinal Chemistry
PrimeKG (Chandak et al., 2023) 10 (e.g., phenotypes) 30 (e.g., disease-phenotype) #N: 129.4 K, #E: 8.1 M Medicinal Chemistry
Gene Ontology [^1] http://geneontology.org/ (Ashburner et al., 2000) 3 (e.g., biological process) 4 (e.g., partOf) #N: 43 K, #E: 7544.6K Bioinformatics
KEGG [^2] https://www.genome.jp/kegg/ (Kanehisa and Goto, 2000) 16 (e.g., pathway) 4 (e.g., partOf) #N: 48.5 M, #E: unknown Bioinformatics
STRING [^3] https://string-db.org/ (Szklarczyk et al., 2023) 1 (e.g., protein) 4 (e.g., interactions) #N: 67.6 M, #E: 20 B Bioinformatics
Cell Ontology [^4] https://www.ebi.ac.uk/ols4/ontologies/cl (Diehl et al., 2016) 1 (i.e., cell type) 2 (e.g, subClassOf) #N: 2.7 K, #E: 15.9 K Bioinformatics
GEFA (Ranjan et al., 2022) 510 (e.g., kinases) 2 (e.g., drug-drug) #N: 0.5 K, #E: 30.1 K Drug Development
Reaction (Li and Chen, 2022) 2 (e.g., reactant & normal) 19 (e.g., reaction paths) #N: 2192.7 K, #E: 932.2 K Drug Development
ASICS (Jeong et al., 2022) 2 (e.g., reactant & product) 1 (e.g., reactions) #N: 1674.9 K, #E: 923.8 K Drug Development
Hetionet (Jeong et al., 2022) 11 (e.g., biological process) 24 (e.g., disease–associates–gene) #N: 47.0 K, #E: 2250.2 K Drug Development
LBD-COVID (Zhang et al., 2021b) 1 (i.e., concept) 1 (i.e., SemMedDB relation) #N: 131.4 K, #E: 1016.1 K Drug Development
GP-KG (Gao et al., 2022) 7 (e.g., drug) 9 (e.g., disease–gene) #N: 61.1 K, #E: 1246.7 K Drug Development
DRKF (Zhang and Che, 2021) 4 (e.g., drug) 43 (e.g., drug-disease) #N: 12.5 K, #E: 165.9 K Drug Development
DDKG (Ghorbanali et al., 2023) 2 (i.e., drug & disease) 1 (e.g., drug-disease) #N: 551, #E: 2.7 K Drug Development
Disease Ontology [^5] https://disease-ontology.org/ (Schriml et al., 2012) 1 (i.e., disease) 2 (e.g., subClassOf) #N: 11.2 K, #E: 8.8 K Clinical Decision Support
DrugBank (Wishart et al., 2018) 4 (e.g., drug, pathway) 4 (e.g., drug-target) #N: 7.4 K, #E: 366.0 K Clinical Decision Support
KnowLife (Ernst et al., 2014) 6 (e.g., genes) 14 (e.g., gene-diseases) #N: 2.9 M, #E: 11.4 M Clinical Decision Support
PharmKG (Zheng et al., 2021) 3 (e.g., diseases) 3 (e.g., chemical-diseases) #N: 7601, #E: 500958 Clinical Decision Support
ROBOKOP [^6] https://robokop.renci.org/ (Bizon et al., 2019) 54 (e.g., genes, drugs) 1064 (e.g., biolink, CHEBI) #N: 8.6M , #E: 130.4 M Clinical Decision Support
iBKH [^7] https://github.com/wcm-wanglab/iBKH (Su et al., 2023) 11 (e.g., anatomy, disease) 18 (e.g., anatomy-gene) #N: 2.4 M, #E: 48.2 M Clinical Decision Support
2.3. HKG Utilization
Model-free Utilization. Various query languages can be used for KGs, such as SPARQL, Cypher, and GraphQL (Wang et al., 2020). These query languages allow users to query healthcare KGs using a standardized syntax, thus enabling users to retrieve, manipulate, and analyze data in a structured and consistent way. More complex applications can be further supported by graph queries. For instance, automatic healthcare question answering can be tackled by Natural Language Question-to-Query (NLQ2Query) approach (Kim et al., 2022), where natural language questions are first translated into executable graph queries and then answered by the query responses. HKGs can also be utilized as an up-to-date and trustworthy augmentation to large language models (LLMs) for many applications. Some pioneering studies (Liu et al., 2022c; Guu et al., 2020; Xu et al., 2023; Shi et al., 2023) show that retrieved knowledge triples can improve the reliability of LLMs in various knowledge-intensive tasks, by addressing the nonsensical or unfaithful generation. Moreover, KGs can be a useful tool for fact-checking (Tchechmedjiev et al., 2019; Vedula and Parthasarathy, 2021; Mayank et al., 2022) as they provide a structured representation of information that can be used to quickly and efficiently verify the accuracy of claims. Researchers have explored the utility of HKGs in identifying ingredient substitutions of food (Shirai et al., 2021), COVID-19 fact-checking (Mengoni and Yang, 2022), etc.
Model-based Utilization. Utilizing HKGs in complex reasoning tasks often involves utilizing machine learning models. HKG embeddings (Yu et al., 2021; Su et al., 2022) have shown great potential to tackle these tasks. In particular, HKG embedding models are a class of machine learning models that aim to learn low-dimensional vector representations, or embeddings, of the entities and relations in a knowledge graph. After obtaining HGK embeddings, they can be plugged into any kind of deep neural network and further fine-tuned toward downstream objectives. On the other hand, symbolic logic models represent another prominent approach for KG reasoning due to their interpretability. More specifically, symbolic reasoning models first mine logical rules from existing knowledge by inductive logic programming (Muggleton, 1992), association rule mining (Galárraga et al., 2013), or Markov logic networks (Kok and Domingos, 2005). These minded rules are used to infer new facts, make logical deductions and answer complex queries. Recently, researchers start to explore combining logical rules into KG embedding to further improve the generalization and performance of HKG reasoning (Alshahrani et al., 2017; Zhu et al., 2022a).
3. Resources
In this section, we compile a detailed resource overview of existing healthcare knowledge graphs to assist researchers and healthcare professionals in constructing and utilizing HKGs, organized in Table 1. We present key attribute information, including HKG name, node types, edge types, statistics, and their applications. Details and external links can be referred to at our public repository [^8] Resource: https://github.com/lujiaying/Awesome-HealthCare-KnowledgeBase.
[…]
L1/.style=draw=black, text width=2.5cm, L2/.style=draw=black, text width=4.2cm, L3/.style=draw=black, text width=4.5cm, L4/.style=draw=black, text width=4.5cm, {forest} for tree=draw,rounded corners,grow’=0,text width=2cm, text centered,edge+=-¿, forked edges, [Application, fill=blue!30, L1 [Basic Science Research, fill=blue!20, L2 [Medical Chemistry, fill=blue!10, L3] [Bioinformatics, fill=blue!10, L3] ] [Pharmaceutical Development, fill=blue!20, L2 [Drug Development, fill=blue!10, L3] [Clinical Trial, fill=blue!10, L3] ] [Clinical Decision Support, fill=blue!20, L2 [Intermediate Steps to Advance Prediction Models, fill=blue!10, L3] [Evidence Generation for Risk Prediction, fill=blue!10, L3] ] [Public Health, fill=blue!20, L2 [Epidemiology, fill=blue!10, L3] [Environmental Health, fill=blue!10, L3] [Policy and Management, fill=blue!10, L3] [Social and Behavioral, fill=blue!10, L3] ] ]
Figure 4. Detailed taxonomy of the application section of healthcare knowledge graphs.
4. Applications
4.1. Basic Science Research
Several previous biological terms can also be considered knowledge graphs, such as ontology (gene ontology, cell ontology, disease ontology), network (gene regulatory network), etc. We use the original biological terms as they are more popular according to historical reasons.
4.1.1. Medicinal Chemistry
Topics related to medicinal chemistry involve drug-drug interactions (DDIs) and drug-target interactions (DTIs), which will be discussed in this section.
Drug-drug interactions (DDIs) refer to changes in the actions, or side effects, of drugs when they are taken at the same time or successively (Giacomini et al., 2007). In general, DDIs are a significant contributor to life-threatening adverse events (Su et al., 2022), and their identification is one of the key tasks in public health and drug development. The existence of diverse datasets on drug-drug interactions (DDIs) and biomedical KGs has enabled the development of machine-learning models that can accurately predict DDIs. Yu et al. (Yu et al., 2021) develop SumGNN, a model that includes a subgraph extraction module to efficiently extract relevant subgraphs from a KG, a self-attention-based summarization scheme to generate reasoning paths within the subgraph, and a multichannel module for integrating knowledge and data, resulting in significantly improved predictions of multi-typed DDIs. Su et al. (Su et al., 2022) propose DDKG, an attention-based KG representation learning framework that involves an encoder-decoder layer to learn the initial embeddings of drug nodes from their attributes in the KG. Karim et al. (Karim et al., 2019) compare various techniques for generating KG embeddings with different settings and conclude that a combined convolutional neural network and LSTM network yields the highest accuracy when predicting drug-drug interactions (DDIs). Dai et al. (Dai et al., 2021) propose a new KG embedding framework by introducing adversarial autoencoders based on Wasserstein distances and Gumbel-Softmax relaxation for DDI tasks. Lin et al. (Lin et al., 2020) develop KGNN that resolves the DDI prediction by capturing drugs and their potential neighborhoods by mining associated relations in KG.
Drug-target interactions (DTIs) is just as important as DDIs (Chen et al., 2016). Machine learning models can leverage knowledge graphs constructed from various types of interactions, such as drug-drug, drug-disease, protein-disease, and protein-protein interactions, to aid in predicting DTIs. For instance, Li et al. (Li et al., 2023) utilize the KG transfer probability matrix to redefine the drug-drug and target-target similarity matrix, thus constructing the final graph adjacent matrix to learn node representations by VGAE and augmenting them by utilizing dual Wasserstein Generative Adversarial Network with gradient penalty. Zhang et al. (Zhang et al., 2021c) propose a new hybrid method for DTI prediction by first constructing DTI-related KGs and then employing graph representation learning model to obtain feature vectors of the KG. Wang et al. (Wang et al., 2022a) construct a knowledge graph of 29,607 positive drug-target pairs by DistMult embedding strategy, and propose a Conv-Conv module to extract features of drug-target pairs. Ye et al. (Ye et al., 2021b) learn a low-dimensional representation for various entities in the KG, and then integrate the multimodal information via neural factorization machine.
4.1.2. Bioinformatics Research
In bioinformatics settings, a knowledge graph is a graph-based representation where nodes are biomedical entities (such as mutations, genes, proteins, metabolites, diseases, and biological pathways), and edges are their relationships (such as associations, interactions, and regulations). Through the integration, researchers can gain a more comprehensive understanding of complex biological processes and diseases.
Multi-Omics Applications: In recent years, the field of multi-omics data analysis has become increasingly important for understanding complex biological systems. With the advancement of high-throughput technologies, researchers can generate large-scale and high-dimensional data from different omics fields, such as genomics, transcriptomics, proteomics, metabolomics, and epigenomics. More KG applications based on multi-omics data integration have emerged, aiming to provide new research methods to uncover the complex relationships between different omics layers and reveal biological systems’ underlying mechanisms.
Knowledge graphs have been used to identify disease-associated mutations, genes, proteins, and metabolites by integrating multi-omics data with existing biological knowledge. This approach has led to the discovery of novel biomarkers and therapeutic targets for various diseases and the interpretation of the functional effects of genetic elements (Zhang et al., 2021a). Quan et al. built a comprehensive multi-relational knowledge graph called AIMedGraph, providing an interpretation of the impact of genetic variants on disease or treatment (Quan et al., [n. d.]). They curated detailed information about diseases, drugs, genetic variants, and the impact of genetic variations on disease development and drug treatment from multiple data resources. The entities integrated into AIMedGraph are connected by evidence-based relations and form a comprehensive gene–variant–disease–drug–trial–reference knowledge network. AIMedGraph uses the Adamic-Adar algorithm to predict new relations between entities based on shared neighbors and their proximity in the knowledge network. GenomicsKG is a knowledge graph to analyze and visualize multi-omics data. GenomicsKG can be used to improve drug development based on clinical genomics correlations and personalized drug customization in the extended version based on interactive relationships. It also provides multi-dimensional visualization, linked functional knowledge graphs, and reporting for clinical genomics.
Single-Cell Analysis: Cells are fundamental and essential units of living organisms. With high-throughput sequencing technologies advancing to measure genomic profiles in a single-cell resolution, cell functions (inside cells) and cell-cell interactions (between cells) are revealed (Linnarsson and Teichmann, 2016). When diving into the functions of cells, gene regulatory mechanisms can be a critical factor. Gene regulatory mechanisms control the expression of genes, which can affect cell differentiation, response to stimuli, disease progression, etc. It can be visualized as gene regulatory networks (GRNs) which depict the interactions between genes and their regulators. Traditional approaches to constructing the GRN are based on gene knockdown or knockout experiments, which can be time-consuming, labor-intensive, and costly. In addition, these experiments are limited to only one or a few genes which neglects gene-gene interactions. In contrast, single-cell sequencing data could provide whole-genome scale measurements in each individual cell with comprehensive information such as gene expression, transcription factors (TF) binding sites, DNA methylation, epigenetic modifications, etc. By mining these publicly available data, it is possible to reveal the underlying and universal GRN biomedical researchers to understand biological processes better. For example, GRNdb provides detailed regulon and TF-target pairs information from different human and mouse tissues under different conditions by analyzing existing sequencing data (Fang et al., 2021). GenomicKB integrates existing datasets and genome annotations and formulates the data into a knowledge graph to emphasize the relationships among genomic entries (Feng et al., 2023). We can foresee that with more data being generated and collected, one generalized GRN (gene regulatory knowledge graph) along with cell-type-specific GRNs can be reconstructed and used to help reveal the underlying mechanisms of gene expression, biological processes, and disease progression.
Besides cell-type-specific GRN, single-cell sequencing data can also be used to infer cell-cell interactions, which play a critical role in understanding cell cycles, cell fate decisions, tissue development, immune response, etc. Among several approaches through which cells can interact, cell-cell communication or cell signaling is of the most interest (Eltzschig et al., 2006) as the physical distance between cells does not limit it. Cells can communicate with each other by sending signaling molecules called ligands and receiving through receptors located on cell surfaces. Several methods have been developed to implicitly measure cell-cell communication by modeling and scoring ligand-receptor interactions (LRIs) from single-cell sequencing data (Efremova et al., 2020; Jin et al., 2021). However, these methods do not utilize the spatial information of cells. Recent advancements in spatial sequencing techniques provide not only genomic measurements but also complementary information on the spatial coordinates of each cell, where the spatial coordinates indicate the probabilities of cells communicating with each other as cells with closer distances tend to interact more. A recent benchmark study on evaluating cell-cell communication inference methods based on spatial distances (Liu et al., 2022a) while another benchmark study also considers other biological factors such as cytokine activities and receptor protein abundance (Dimitrov et al., 2022). With the availability of spatial transcriptomics data, many methods have been developed to leverage colocalization information to infer LRIs. Although many LRIs databases (Efremova et al., 2020; Shao et al., 2021) have been constructed and applied to infer cell-cell communication, only recently, SpaTalk (Shao et al., 2022) integrates CellTalkDB, Kyoto Encyclopedia of Genes and Genomes (KEGG) pathways, Reactome and TFs from AnimalTFDB to construct a ligand-receptor-target knowledge graph to help improve the inference of cell-cell communication. The success of SpaTalk provides an unprecedented view of integrating existing networks into a knowledge graph and shows that correctly utilizing the knowledge graph could provide helpful information in cell-cell communication tasks.
4.2. Pharmaceutical Research Development
4.2.1. Drug Development
Drug development is the process of identifying novel chemical compounds that can effectively treat or alleviate human diseases. Before the drug can be designated as a final product for clinical use, several critical steps need to be undertaken from the initial target identification, chemical synthesis and clinical trials. The whole process typically spans over a decade and involves expenses of approximately one billion dollars (Adams and Brantner, 2006), yet it is characterized by a low success rate for clinical approval (DiMasi et al., 2015).
With advancements in chemistry and the availability of vast chemical libraries, it is possible to reduce the costs of drug development by extracting and integrating valuable insights from existing data and using computer algorithms to accelerate the drug design process (Du et al., 2021; Zeng et al., 2022; Wang et al., 2022c; Pan et al., 2022b; Wang et al., 2022b; Nicolaou et al., 2007; Nicolaou and Pattichis, 2006). Despite the growing trend of computer-assisted drug discovery, a key question remains regarding how to effectively integrate data and extract valuable insights from the vast chemical dataset.
To approach this question, knowledge graphs (KGs) have been employed for drug discovery due to their various advantages (Zeng et al., 2022): (1) In contrast to traditional methods that capture only one type of relationship, KGs are capable of providing heterogeneous information that includes diverse entities (e.g., scaffolds, proteins, and genes); (2) In addition, KGs are capable of handling multiple types of relationships between various types of entities, such as drug-target pairs; and (3) KGs can provide unstructured semantic relationships between entities. In such graphs, entities are represented as nodes while their relationships are represented as edges, by which complex relations in biochemical systems can be easily handled.
In general, the field of drug development encompasses two main areas: drug design and drug repurposing (also known as drug repositioning). Drug design is the process of creating novel and diverse drug molecules with desirable pharmaceutical properties (Jing et al., 2018; Fu et al., 2021; Fu and Sun, 2022), whereas drug repurposing identifies new uses for existing approved drugs that were originally developed for a different indication (Pan et al., 2022a; Huang et al., 2020).
Drug Design. KGs are widely employed in drug design, particularly in generating novel molecules that are promising drug candidates for various diseases (Ranjan et al., 2022; Li and Chen, 2022). Ranjan et al. (Ranjan et al., 2022) utilize Gated Graph Neural Network (GGNN) to generate novel molecules that target the coronavirus (i.e., SARS-CoV-2) (Hasöksüz et al., 2020) and integrate KGs into their approach to reduce the search space. Specifically, KGs were leveraged to discard non-binding molecules before inputting them into the Early Fusion model, thus optimizing the efficiency of the drug design process. In addition to employing deep learning for direct structure design, KGs are also utilized in the analysis of chemical synthesis. Quantitative estimation of molecular synthetic accessibility is critical in prioritizing the molecules generated from generative models. For instance, Li et al. (Li and Chen, 2022) utilize reaction KGs to construct classification models for compound synthetic accessibility. By leveraging KGs that capture information about reactions, including reaction types, substrates, and reaction conditions, they can train machine learning models that could predict the synthetic accessibility of compounds. Jeong et al. (Jeong et al., 2022) introduce an intelligent system that integrates generative exploration and exploitation of reaction knowledge base to support synthetic path design.
Drug Repurposing. Compared to drug design, KGs are more commonly utilized to expedite the drug re-purposing process (Zhu et al., 2020; MacLean, 2021; Himmelstein et al., 2017; Zhang et al., 2021b; Gao et al., 2022; Xu et al., 2019; Zhang and Che, 2021; Ghorbanali et al., 2023). Many applications on drug re-purposing that utilize KGs are primarily focused on link prediction tasks (MacLean, 2021). To re-purpose promising drug candidates for new indications, many methods employ predictive models that focus on predicting drug-treats-disease relationships within pharmacological knowledge graphs KGs. Himmelstein et al. (Himmelstein et al., 2017) use a degree-normalized pathway model on the hetionet KG, which includes genes, diseases, tissues, pathophysiologies, and multimodal edges, to identify potentially repurposable drugs for epilepsy. Xu et al. (Xu et al., 2019) develop a multi-path random walk model on a network that incorporates gene-phenotype associations, protein-protein interactions, and phenotypic similarities for training and prediction purposes. Zhang et al. (Zhang et al., 2021b) introduce an integrative and literature-based discovery model for identifying potential drug candidates from COVID-19-focused research literature, including PubMed and other relevant sources. Gao et al. (Gao et al., 2022) construct a knowledge graph (KG) by integrating multiple genotypic and phenotypic databases. They then learn low-dimensional representations of the KG and utilize these representations to infer new drug-disease interactions, providing insights into potential drug repurposing opportunities. Zhang and Che (Zhang and Che, 2021) introduce a model for drug re-purposing in Parkinson’s disease that leverages a local medical knowledge base incorporating accurate knowledge along with medical literature containing novel information. Ghorbanali et al. (Ghorbanali et al., 2023) present the DrugRep-KG method, which utilizes a KG embedding approach for representing drugs and diseases in the process of drug repurposing.
4.2.2. Clinical Trial
The major goal of clinical trials is to assess the safety and effectiveness of drug molecules on human bodies. A novel drug molecule needs to pass three phases of clinical trials before it is approved by the Food and Drug Administration (FDA) and enters the drug market. The whole process is prohibitively time-consuming and expensive, costing 7-11 years and two billion dollars on average (Martin et al., 2017).
Clinical Trial Optimization targets identifying eligible patients for clinical trials based on their medical history and health conditions (Rivera et al., 2020; He et al., 2020). Recently, with massive electronic health records (EHR) data and trial eligibility criteria (EC), data-driven methods have been studied to automatically assign appropriate patients for clinical trials (Yuan et al., 2019; Tseo et al., 2020; Liu et al., 2021). However, it is often hard to fully capture and represent the complex knowledge present in unstructured ECs and EHR data, as ECs may only provide general disease concepts. In contrast, patient EHR data contain more specific medical codes to represent patient conditions. To better capture the interactions among different medical concepts from EHR records and ECs, Gao et al. (Gao et al., 2020) enhance patient records with hierarchical taxonomies to align medical concepts of varying granularity between EHR codes and ECs. Besides, Fu et al. (Fu et al., 2022) leverage additional knowledge-embedding modules along with drug pharmacokinetic and historical trial data to improve the patient trial optimization process, and Wang et al. (Wang et al., 2023) leverage the knowledge graphs to learn static trial embedding and further designed meta-learning module to generalize well over the imbalanced clinical trial distribution.
4.3. Clinical Decision Support
Nowadays, abundant Electronic Health Record (EHR) data enables better computational models for accurate diagnoses and treatments. EHR contains essential patient information such as disease diagnoses, prescribed medications, and test results. Due to this valuable information, EHRs are extensively utilized to identify patterns in patient health and assist healthcare providers in making informed clinical decisions.
However, the sparsity of EHR data typically allows for only a small fraction of medical codes to be learned effectively, thereby restricting the ability of deep learning approaches. To overcome this drawback, knowledge graphs have been applied to incorporate prior medical knowledge for these deep learning models, which augment the representation of medical codes to better support the downstream prediction tasks.
4.3.1. Intermediate Steps to Advance Prediction Models
ICD Coding aims to extract diagnosis and procedure codes from clinical notes, which often consisted of raw text (Mullenbach et al., 2018; Zhang et al., 2020; Vu et al., 2021; Dong et al., 2022). It is often challenging, as the size of the candidate target codes can be large and the distribution of the codes is often long-tailed (Kim and Ganapathi, 2021). To overcome this, Xie et al. (Xie et al., 2019) and Cao et al. (Cao et al., 2020a) propose to leverage knowledge graphs as distant supervision (Min et al., 2013; Liang et al., 2020), and inject the label information via structured knowledge graph propagation by leveraging graph convolution networks (Kipf and Welling, 2016) to learn the correlations among medical codes. Besides, Lu et al. (Lu et al., 2020) propose to leverage knowledge graphs and the co-occurrence graph among clinical nodes simultaneously with a knowledge aggregation module to boost the ICD coding performance. Ren et al. (Ren et al., 2022) design a learning curriculum based on the hierarchical structure of the code to address the highly imbalanced label distribution issue and balance between frequent and rare labels. Overall, injecting additional knowledge with graph neural networks offers a way to mitigate the imbalanced label distribution issue and thus better.
Entity and Relation Extraction from Health Records. Health records contain rich unstructured or semi-structured data, making it difficult for clinicians to access and analyze relevant information. Entity and relation extraction helps convert this unstructured text into structured data that can be more easily processed, understood, and utilized. Specifically, entity extraction aims to identify entity mentions from clinical-free texts. There are two key steps for entity extraction, i.e., named entity recognition (NER) and disambiguation (NED). By leveraging additional knowledge graphs, Varma et al. (Varma et al., 2021) transfer structural knowledge from the knowledge base to the medical domain, improving rare entities’ disambiguation accuracy. Yuan et al. (Yuan et al., 2023) inject additional knowledge from the knowledge graphs for entity linking and proposed two additional strategies, namely Post-pruning and Thresholding, to improve the efficiency and remove the effect of unlinkable entity mention. Fries et al. (Fries et al., 2021) leverage clinical ontologies to provide weak supervision sources to create additional training data for clinical entity disambiguation. Besides, relation extraction aims to identify and classify relationships between entities in unstructured text, which facilitates understanding complex biological processes, drug interactions, and disease mechanisms. To incorporate the external knowledge graph, several works (Fei et al., 2021; Roy and Pan, 2021) proposed additional post-training steps to align the language models with biomedical knowledge. Hong et al. (Hong et al., 2021) construct embeddings for a wide range of codified concepts from EHRs to identify relevant features related to a disease of interest, and Lin et al. (Lin et al., 2022) design a co-training scheme to jointly learn from text and knowledge graphs for extracting and classifying disease-disease relations. In summary, fusing knowledge graphs with language models can flexibly accommodate missing data types and bring additional performance gains, especially for those rare entities and relations.
4.3.2. Evidence Generation for Risk Prediction Models
Disease Prediction aims to predict the potential diseases of a given patient with his past clinical records. To assist the diagnosis with additional knowledge, GRAM (Choi et al., 2017) and KAME (Ma et al., 2018) utilize a medical ontology (Dubberke et al., 2006) where the leaf nodes are the medical codes found in EHR data, and their ancestors are more general categories. By incorporating information from medical ontologies into deep learning models via neural attention, these approaches learn better embeddings for different medical concepts to alleviate the data scarcity bottleneck. (Yin et al., 2019; Zhang et al., 2019) further consider the domain-specific knowledge graph KnowLife (Ernst et al., 2015) to enrich the embeddings of medical entities with their neighbors on the knowledge graph. These approaches mainly directly update the embeddings of different concepts to improve the feature learning, but may be at the risk of ignoring the high-level order information from the knowledge graph. To tackle this drawback, Ye et al. (Ye et al., 2021a) explicitly exploit paths in KG from the observed symptoms to the target disease to model the personalized information for diverse patients with a relational-guided attention mechanism. Xu et al. (Xu et al., 2021) design a self-supervised learning approach to pre-train a graph attention network for learning the embedding of medical concepts and completing the knowledge graph simultaneously. These approaches better harness the structure information, and often lead to better performance than the pure embedding-based knowledge integration techniques.
Treatment Recommendation aims to recommend personalized medications to patients based on their individual health conditions, which can help physicians select the most effective medications for their patients, and improve treatment outcomes (Zhang et al., 2017; Bhoi et al., 2021; Shang et al., 2019b). To effectively exploit external knowledge, Shang et al. (Shang et al., 2019a) use drug ontologies to design additional pretraining loss and directly improve the representation of drugs, and several studies (Wu et al., 2022a; Tan et al., 2022) attempt to extract the additional drug interaction graphs to model the negative side effects of specific drug pairs and reduce the possibility of recommending negative drug-drug interaction combinations. Besides, Wu et al. (Wu et al., 2022b) leveraged ontologies to improve the drug representations, and facilitates drug recommendation under a more challenging few-shot setting.
4.4. Public Health
Public Health research can significantly benefit from HKGs. Knowledge graphs can help organize, structure, and formalize extensive information from diverse and heterogeneous sources. This allows researchers to analyze data, reason about factors, and make decisions on a larger scale.
Epidemiology. The field of epidemiology has seen an increase in the use of knowledge graphs to analyze and understand the spread of diseases. A study conducted by Gao et al. (Gao et al., 2023) analyzes the research hotspots and development trends of wastewater-based epidemiology (WBE) using knowledge graphs constructed from nearly 900 papers. Domingo-Fernández et al. (Domingo-Fernández et al., 2021) create the COVID-19 Knowledge Graph, a comprehensive cause-and-effect network constructed from the scientific literature on the coronavirus. Additionally, Turki et al. (Turki et al., 2022) use knowledge graphs to assess and validate the portion of Wikidata related to COVID-19 epidemiology using an automatable task set. Pressat Laffouilhère et al. (Pressat Laffouilhère et al., 2022) develop OntoBioStat, a domain ontology related to covariate selection and bias in biostatistics, which can help interpret significant statistical associations between variables.
Environmental Health. Fecho et al. (Fecho et al., 2021) develop ROBOKOP, a biomedical knowledge graph-based system, to validate associations between workplace chemical exposures and immune-mediated diseases. Wolffe et al. (Wolffe et al., 2020) propose using knowledge graphs in systematic evidence mapping in environmental health. This approach overcomes the limitations of rigid data tables by offering a more suitable model for handling the highly connected and complex nature of environmental health data.
Health Policy and Management. Wu et al. (Wu, 2021) have analyzed the COVID-19 epidemic situation using a knowledge graph of patient activity. This method enables in-depth study of the transmission process, analysis of key nodes, and tracing of activity tracks. Meanwhile, Yu et al. (Yu et al., 2022) develop a chronic management system, which combines knowledge graphs and big data to optimize the management of chronic diseases in children. This system enhances treatment and resource utilization while conforming to the requirements of the Chronic Care Model.
Social and Behavioral Health. There have been several interesting research studies. Cao et al. (Cao et al., 2020b) build a high-level suicide-oriented knowledge graph combined with deep neural networks for detecting suicidal ideation on social media platforms. Also, Liu et al. (Liu et al., 2020) conduct a bibliometric analysis of driver behavior research. Additionally, Wang et al. (Wang et al., 2022d) create an analysis framework for interpreting causal associations in emotional logic. They introduce a knowledge graph into appraisal theories, improving human emotional inference.
5. Promise and Outlook
The potential impact of comprehensive and fine-grained HKGs on biomedical research and clinical practice is significant. By integrating vast amounts of biomedical knowledge from multiple domains, HKGs can facilitate the discovery of new disease mechanisms and the identification of novel drug targets. They also help to enable personalized medicine by identifying patient subgroups with shared disease mechanisms. The recent success of large language models (a.k.a. foundation models) such as ChatGPT offers promising opportunities in capturing such semantics from the biomedical context (Agrawal et al., 2022; Singhal et al., 2022; Nath et al., 2022; Moor et al., 2023), enabling the construction of unprecedentedly comprehensive HKGs. In turn, HKGs also help improve LLMs by providing accurate and contextualized knowledge to regularize the generated content. This is particularly useful in evaluating LLMs in biomedical applications and addressing the problem of hallucination in critical areas.
6. Conclusion
Healthcare knowledge graphs have emerged as a promising approach for capturing and organizing medical knowledge in a structured and interpretable way. This comprehensive review paper provides an overview of the current state of HKGs, including their construction, utilization models, and applications in healthcare. Furthermore, the paper discusses the potential future developments of HKGs. In conclusion, HKGs have played a significant role in advancing health research. With the advent of LLMs, there are even more opportunities to combine HKGs and LLMs to reduce the generation of false or unreliable content. We hope that our comprehensive review of this field offers a helpful perspective for future reference.

### v5

@@
-A Review on Knowledge Graphs for Healthcare: Resources, Applications, and Promises
-Abstract.
-Healthcare knowledge graphs (HKGs) are valuable tools for organizing biomedical concepts and their relationships.
-The recent advance of large language models (LLMs) has paved the way for building more comprehensive and accurate HKGs.
-This, in turn, can improve the reliability and evaluation of LLMs.
-However, the challenges and opportunities of HKGs are not fully understood, highlighting the need for detailed reviews.
-This work provides the first comprehensive review of HKGs, summarizing the pipeline and key techniques for HKG construction and successful HKG utilization in various health-related applications.
-Lastly, we highlight the opportunities for HKGs in the era of LLMs.
-knowledge graph, healthcare, language models, multimodality, interpretability, trustworthy AI
-$c$s: Applied computing Health informaticsccs: Applied computing Health care information systemsccs: Applied computing Bioinformaticsccs: Information systems Graph-based database modelsccs: Computing methodologies Knowledge representation and reasoning
-1. Introduction
-Figure 1. The content overview of this review on healthcare knowledge graph.
-A knowledge graph (KG) is a data structure that captures the relationships between different entities and their attributes (Ji et al., 2021; Nicholson and Greene, 2020).
-KG models and integrates data from various sources, including structured and unstructured data, and has been studied to support a wide range of applications such as search engines (Wang et al., 2019b), recommendation systems (Wang et al., 2019a; Zhou et al., 2020), and question answering (Lin et al., 2019; Yasunaga et al., 2021; Yan et al., 2021; Kan et al., 2021).
+
+Abstract
+Objective: This comprehensive review aims to provide an overview of the current state of Healthcare Knowledge Graphs (HKGs), including their construction, utilization models, and applications across various healthcare and biomedical research domains.
+Methods: We thoroughly analyzed existing literature on HKGs, covering their construction methodologies, utilization techniques, and applications in basic science research, pharmaceutical research and development, clinical decision support, and public health.
+The review encompasses both model-free and model-based utilization approaches and the integration of HKGs with large language models (LLMs).
+Results: We searched Google Scholar for relevant papers on HKGs and classified them into the following topics: HKG construction, HKG utilization, and their downstream applications in various domains.
+We also discussed their special challenges and the promise for future work.
+Discussion: The review highlights the potential of HKGs to significantly impact biomedical research and clinical practice by integrating vast amounts of biomedical knowledge from multiple domains.
+The synergy between HKGs and LLMs offers promising opportunities for constructing more comprehensive knowledge graphs and improving the accuracy of healthcare applications.
+Conclusions: HKGs have emerged as a powerful tool for structuring medical knowledge, with broad applications across biomedical research, clinical decision-making, and public health.
+This survey serves as a roadmap for future research and development in the field of HKGs, highlighting the potential of combining knowledge graphs with advanced machine learning models for healthcare transformation.
+1 Introduction
+A knowledge graph (KG) is a data structure that captures the relationships between different entities and their attributes 1; 2.
+KG models and integrates data from various sources, including structured and unstructured data, and has been studied to support a wide range of applications such as search engines 3, recommendation systems 4; 5, and question answering 6; 7; 8; 9.
 Healthcare Knowledge Graph (HKG) facilitates an interpretable representation of medical concepts, e.g., drugs and disease, as well as the relations among those medical concepts.
-This data structure enables the connection of contexts and enhances clinical research and decision-making (Santos et al., 2022; Chandak et al., 2023).
-On the data side, HKG is usually built on complex medical systems such as electronic health records, medical literature, clinical guidelines, and patient-generated data (Bouayad et al., 2017; Rajkomar et al., 2018).
-However, these data resources are often heterogeneous and distributed, which makes it challenging to integrate and analyze them effectively (Mehta and Pandit, 2018).
-This data heterogeneity can also lead to incomplete or inconsistent data representations, limiting their usefulness for downstream healthcare tasks (Dash et al., 2019).
-Additionally, the current use of domain-specific knowledge graphs may result in limited coverage and granularity of the knowledge captured across different levels.
+This data structure enables the connection of contexts and enhances clinical research and decision-making 10; 11.
+On the data side, HKG is usually built on complex medical systems such as electronic health records, medical literature, clinical guidelines, and patient-generated data 12; 13.
+However, these data resources are often heterogeneous and distributed, which makes it challenging to integrate and analyze them effectively 14.
+This data heterogeneity can also lead to incomplete or inconsistent data representations, limiting their usefulness for downstream healthcare tasks 15.
+Additionally, the current use of domain-specific KGs may result in limited coverage and granularity of the knowledge captured across different levels.
 This hinders identifying correlations and relationships between medical concepts from multiple domains.
@@
 On the modeling side, the construction of HKGs can be done either from scratch or by integrating existing dataset resources.
-Many crucial steps, such as entity and relation extraction, can be optimized with natural language processing tools and algorithms.
-Recently, there have been significant advancements in general domain knowledge extraction, thanks to pre-trained large language models (LLMs) such as BERT (Devlin et al., 2019), GPT Series (Brown et al., 2020), and others.
+Many crucial steps, such as entity and relation extraction, can be done with NLP tools and algorithms.
+Recently, there have been significant advancements in general domain knowledge extraction, thanks to pre-trained large language models (LLMs) such as BERT 16, GPT Series 17, and others.
 These models revolutionize the field and make it possible to integrate heterogeneous medical data from various sources effectively.
-The use of pre-trained models has also led to the development of more accurate and comprehensive medical ontologies and taxonomies (Zhang et al., 2021d; Yu et al., 2020; Wang et al., 2021; Zeng et al., 2021; Xu et al., 2020).
+The use of pre-trained models has also led to the development of more accurate and comprehensive medical ontologies and taxonomies 18; 19; 20; 21; 22.
 This allows for the evaluation of generated contents from LLMs and reduces LLM hallucination.
-A comprehensive healthcare knowledge graph has the potential to contribute to health research across various levels (Gyrard et al., 2018; Santos et al., 2022; Li et al., 2020).
-At the micro-scientific level, HKGs can help researchers identify new phenotypic and genotypic correlations and understand the underlying mechanisms of disease (Hassani-Pak and Rawlings, 2017), leading to more targeted and effective treatments (Seneviratne et al., 2021; Chandak et al., 2023).
-At the clinical care level, HKGs can be used to develop clinical decision support systems that provide clinicians with relevant information, improving clinical workflows and patient outcomes (Eberhardt et al., 2012; Castaneda et al., 2015).
+A comprehensive HKG has the potential to contribute to health research across various levels 23; 10; 24.
+At the micro-scientific level, HKGs can help researchers identify new phenotypic and genotypic correlations and understand the underlying mechanisms of disease 25, leading to more targeted and effective treatments 26; 11.
+At the clinical care level, HKGs can be used to develop clinical decision support systems that provide clinicians with relevant information, improving clinical workflows and patient outcomes 27; 28.
 Therefore, a thorough review of existing literature on HKGs becomes an essential roadmap and invaluable resource to drive transformative advancements in the field.
-This review is the first comprehensive overview of HKGs, covering contents shown in Figure 1.
-Specifically, in Section 2, we delve into the construction pipelines of HKGs, including both building from scratch and integration approaches, and highlight the key techniques employed in HKG construction.
-Additionally, we explore two standard utilization methods of HKGs, namely model-free and model-based approaches.
-In Section 3, we compile a comprehensive summary of existing HKG resources with their scopes and applications.
-Furthermore, Section 4 investigates the literature on mainstream health applications, offering an in-depth overview of the diverse use cases of HKGs in healthcare.
-Finally, we address promising research opportunities in the era of LLMs in Section 5.
-2. Backgrounds
-2.1. Knowledge Graphs for Healthcare (HKG)
-Definition.
-A healthcare knowledge graph (HKG) is a domain-specific knowledge graph designed to capture medical concepts such as drugs, diseases, genes, phenotypes, and so on, and their relationships in a structured and semantic way.
-Terminology.
-Our focus is on healthcare knowledge graphs (HKGs), which are structured, semantic representations of medical concepts and relationships.
-We also include ontologies and knowledge bases, which are commonly used in constructing HKGs.
-Ontologies define a set of concepts and categories in a domain, as well as the relationships between them, while knowledge bases store factual information about entities and their attributes.
+Statement of Significance: While prior reviews have focused on KGs tailored for specific healthcare tasks such as drug discovery 29, electronic health record (EHR) predictions 30, and ontologies 31, they often lack a holistic perspective across diverse healthcare domains.
+General surveys on KG construction 1 provide foundational insights but do not delve into the unique challenges and requirements of the healthcare context.
+Furthermore, some studies 32 concentrate solely on KG construction without exploring the critical dimension of knowledge utilization in healthcare applications.
+This work bridges these gaps by offering a comprehensive review that not only synthesizes knowledge graph construction techniques but also emphasizes their practical applications across various healthcare domains, providing a roadmap for future innovation.
+Figure 1: The content overview of this review on healthcare knowledge graphs.
+Literature Selection.
+We conducted a systematic literature review following the PRISMA (Preferred Reporting Items for Systematic Reviews and Meta-Analyses) guidelines to identify and include studies relevant to the development and application of HKGs.
+Searches were conducted across multiple academic databases, including Google Scholar, PubMed, IEEE Xplore, ACM Digital Library, ACL Anthology, and Web of Science.
+Search terms were crafted to capture the scope of HKG research comprehensively.
+These terms included combinations of the following keywords using Boolean operators (AND, OR, NOT):
+Healthcare Knowledge Graphs: “healthcare knowledge graph” “HKG” “Ontology” “medical knowledge graph”.
+Applications: “clinical decision support” “drug repurposing” “pharmaceutical development” “clinical trial matching” “bioinformatics”.
+Construction Techniques: “ontology integration” “entity resolution” “relation extraction” “knowledge graph embeddings”.
+Integration with AI: “large language models” “machine learning” “natural language processing”.
+Reference lists of key articles were manually searched to identify additional relevant studies.
+The search timeframe spanned from 2010 to 2024 with 362 articles in total.
+We then removed duplicate articles based on their titles and authors.
+After deduplication, we examined the remaining articles’ titles and abstracts for topic relevance.
+Finally, for each section, we conducted a detailed examination of the methodology described in each paper to ensure relevance, excluding those deemed irrelevant.
+After the filtering process, a total of 175 papers were retained.
+Figure 2: The PRISMA diagram illustrating the literature search process.
+2 Backgrounds
+Healthcare Knowledge Graphs (HKGs).
+We focus on HKG, which is a structured representation of knowledge that captures entities (e.g., organizations, diseases, medications) and the relationships between them.
+KGs typically employ graph-based structures, where entities are represented as nodes and relationships as edges, and may use attributes or properties for additional contexts.
+These graphs can be enriched with contextual information and metadata, enabling complex relationships and data interconnectivity.
+Besides, we also include ontologies and knowledge bases, which are commonly used in constructing HKGs.
+An ontology is a formal model of a domain’s concepts, properties, and relationships, typically using a hierarchical or taxonomical structure, with an emphasis on semantic relation, and interoperability of different healthcare concepts.
+In contrast, KGs integrate data from heterogeneous sources, scale dynamically, and are well-suited for data-driven insights and probabilistic reasoning.
+By leveraging both ontologies and KGs, HKGs can support robust knowledge discovery and reasoning for healthcare applications.
+Furthermore, combining these resources with LLMs can enhance decision support by grounding LLM outputs in structured, contextualized knowledge and reduce the risk of hallucination.
 By covering these categories of terminology, we provide a comprehensive overview of the different types of resources available for organizing and representing medical knowledge in a structured and semantically rich manner.
@@
 L1/.style=draw=black, text width=3cm, L2/.style=draw=black, text width=4cm, L3/.style=draw=black, text width=5cm, L4/.style=draw=black, text width=5cm, {forest} for tree=draw,rounded corners,grow’=0,text width=2cm, text centered,edge+=-¿, forked edges, [Background, fill=orange!30, L1 [HKG Definition, fill=orange!20, L2] [HKG Construction, fill=orange!20, L2 [Construct from Scratch, fill=orange!10, L3] [Construct by Integration, fill=orange!10, L3] ] [HKG Utilization, fill=orange!20, L2 [Model-free Utilization, fill=orange!10, L3] [Model-based Utilization, fill=orange!10, L3] ] ]
-Figure 2. Detailed taxonomy of the background section of healthcare knowledge graphs.
-2.2. HKG Construction
-Healthcare knowledge graphs can be constructed from scratch or through the integration of existing data resources.
+Figure 3: Detailed taxonomy of the background section of healthcare knowledge graphs.
+2.1 HKG Construction
 Develop schemaCollect dataExtract knowledge triplesNormalize entities and relationsInfer missing linksUpdate and validate
-Figure 3. The pipeline of constructing HKGs from scratch.
+Figure 4: The pipeline of constructing HKGs from scratch.
 Constructing HKGs from Scratch. A multi-step pipeline, as in Figure 4, is used to construct HKGs from scratch.
-  1. (1)
+  1. 1.
     The first step is to identify the scope and objectives.
-In most cases, researchers develop a schema (Guarino et al., 2009; Blagec et al., 2022) or use existing schemas (Guarino et al., 2009; Ashburner et al., 2000; Schriml et al., 2012; Bard et al., 2005) to serve as the formal and explicit specification of a domain, thus ensuring consistent, coherent, and aligned domain knowledge.
+Researchers typically develop a schema 33; 34 or use existing schemas 33; 35; 36; 37 to clearly specify the domain, ensuring consistent and aligned knowledge.
 Unlike the general domain KG, utilizing schemas is a common practice in HKG construction.
-  2. (2)
+  2. 2.
     Secondly, researchers gather data from various sources, including medical literature, clinical trials, and patient-generated data.
 It is essential to ensure the quality and consistency of the data and to remove identifiable information for patient privacy.
-  3. (3)
-    The third step is to extract and transform the data into a structured format.
-This step involves identifying medical entities and creating relationships between them via specialized biomedical Natural Language Processing (NLP) tools (Song et al., 2021; Xing et al., 2020; Hahn and Oleynik, 2020).
-  4. (4)
-    Next, researchers map the entities and relationships to the chosen ontologies with the help of thesauruses (Bodenreider, 2004) or terminologies (Donnelly et al., 2006; Hirsch et al., 2016).
-This step ensures that the knowledge graph is interoperable with other healthcare systems and facilitates data integration.
-  5. (5)
-    Until now, an initial KG has been built.
-The next step is to populate the KG to infer missing links between entities.
-This inference can be done using graph databases (Wang et al., 2020) or link prediction models (Bordes et al., 2013; Lu and Yang, 2022).
-  6. (6)
+  3. 3.
+    The third step involves transforming the data into a structured format.
+This includes identifying medical entities and creating relationships between them via specialized biomedical NLP tools 38; 39; 40.
+  4. 4.
+    Entities and relationships are then mapped to selected ontologies using thesauruses 41 or terminologies 42; 43 to ensure the KG is compatible with other healthcare systems and supports data integration.
+  5. 5.
+    Once the initial KG is built, the next step is to fill in missing links between entities using graph databases 44 or link prediction models 45; 46.
+  6. 6.
     The final step is to continuously update and validate the KG to ensure accuracy and relevance.
@@
 Considering significant efforts have been paid to construct and curate HKGs from scratch, it is promising to integrate these data resources to avoid repetitive work.
-Healthcare KG integration (also called Healthcare KG fusion) refers to the processing of merging two or more HKGs into a single, more comprehensive graph (Himmelstein et al., 2017; Su et al., 2023; Youn et al., 2022).
+HKG integration (also called HKG fusion) refers to the processing of merging two or more HKGs into a single, more comprehensive graph 47; 48; 49.
 The integration process is challenging because different HKGs may use different terminologies, schemas, or data formats.
-To address these challenges, researchers have developed various techniques and algorithms for knowledge graph fusion, including ontology matching (Faria et al., 2014; He et al., 2022), schema alignment (Suchanek et al., 2011; Maaroufi et al., 2014), entity resolution (Bachman et al., 2018; Hu et al., 2021), and conflicts resolution (Ma et al., 2023).
+To address these challenges, researchers have developed various techniques and algorithms for KG fusion, including ontology matching 50; 51, schema alignment 52; 53, entity resolution 54; 55, and conflicts resolution 56.
 These methods aim to identify and reconcile the differences between KGs.
@@
 Traditionally, each step of HKG construction involves one specially designated model.
-For instance, Hidden Markov Models and Recurrent Neural Networks are widely used for healthcare named entity recognition, relation extraction, and other sequence tagging tasks, while Translational Models and Graph Neural Networks are used for HKG completion and conflicts resolution tasks.
-Recently, large language models (LLMs) have shown great utility to serve as a uniform tool for constructing KGs (Ye et al., 2022).
-Several key steps of constructing KGs, such as named entity recognition (Liang et al., 2020; Chen et al., 2023; Huang et al., 2022; Liu et al., 2022b), relation extraction (Zhuang et al., 2022; Lu et al., 2022; Yang et al., 2022), entity linking (De Cao et al., [n.
-d.]; Mrini et al., 2022; Cho et al., 2022), and KG completion (geng2022disentangled; Saxena et al., 2022; Xie et al., 2022; Shen et al., 2022), have been successfully tackled by these large foundation models.
-Early explorations of construction HKG with large foundation models show that healthcare entity normalization (Zhang et al., 2023; Agrawal et al., 2022), healthcare entity recognition (Fries et al., 2021; Hu et al., 2023), healthcare entity linking (Zhu et al., 2022b), and healthcare knowledge fusion (Lu et al., 2023) can also be performed, without extensive training on expensive healthcare annotated corpus.
-On the other hand, researchers start to construct KGs under the open-world assumption (Shi and Weninger, 2018; Das et al., 2020; Niu et al., 2021; Li et al., 2022; Lu and Yang, 2022), thus getting rid of the dependency on pre-defined schemas and exhaustive entity&relation normalization.
-Although open-world KGs greatly increase the coverage, ensuring the quality of extracted knowledge is still an open research challenge, especially for explainable and trustworthy HKGs.
-Table 1. Resource of existing healthcare knowledge graph (HKG).
+For instance, Hidden Markov Models and Recurrent Neural Networks are widely used for healthcare named entity recognition, relation extraction tasks, while Translational Models and Graph Neural Networks are used for HKG completion and conflict resolution tasks.
+Recently, LLMs have shown great utility to serve as a uniform tool for constructing KGs 57.
+Several key steps of constructing KGs, such as named entity recognition 58; 59; 60; 61, relation extraction 62; 63; 64, entity linking 65; 66; 67, and KG completion 68; 69; 70, have been successfully tackled by these large foundation models.
+Early explorations of construction HKG with large foundation models show that healthcare entity normalization 71; 72, healthcare entity recognition 73; 74, healthcare entity linking 75, and healthcare knowledge fusion 76 can also be performed, without extensive training on healthcare corpus.
+On the other hand, researchers start to construct KGs under the open-world assumption 77; 78; 79; 80; 46, thus getting rid of the dependency on pre-defined schemas and exhaustive entity&relation normalization.
+Table 1: Resources of existing HKGs.
 Name Node Types Edge Types Statistic Application
-HetioNet (Himmelstein and Baranzini, 2015) 11 (e.g., drug, disease) 24 (e.g., drug-disease) #N: 47.0 K, #E: 2.3 M Medicinal Chemistry
-DrKG (Ioannidis et al., 2020) 13 (e.g., disease, gene) 107 (e.g., disease-gene) #N: 97 K, #E: 5.8 M Medicinal Chemistry
-PrimeKG (Chandak et al., 2023) 10 (e.g., phenotypes) 30 (e.g., disease-phenotype) #N: 129.4 K, #E: 8.1 M Medicinal Chemistry
-Gene Ontology [^1] http://geneontology.org/ (Ashburner et al., 2000) 3 (e.g., biological process) 4 (e.g., partOf) #N: 43 K, #E: 7544.6K Bioinformatics
-KEGG [^2] https://www.genome.jp/kegg/ (Kanehisa and Goto, 2000) 16 (e.g., pathway) 4 (e.g., partOf) #N: 48.5 M, #E: unknown Bioinformatics
-STRING [^3] https://string-db.org/ (Szklarczyk et al., 2023) 1 (e.g., protein) 4 (e.g., interactions) #N: 67.6 M, #E: 20 B Bioinformatics
-Cell Ontology [^4] https://www.ebi.ac.uk/ols4/ontologies/cl (Diehl et al., 2016) 1 (i.e., cell type) 2 (e.g, subClassOf) #N: 2.7 K, #E: 15.9 K Bioinformatics
-GEFA (Ranjan et al., 2022) 510 (e.g., kinases) 2 (e.g., drug-drug) #N: 0.5 K, #E: 30.1 K Drug Development
-Reaction (Li and Chen, 2022) 2 (e.g., reactant & normal) 19 (e.g., reaction paths) #N: 2192.7 K, #E: 932.2 K Drug Development
-ASICS (Jeong et al., 2022) 2 (e.g., reactant & product) 1 (e.g., reactions) #N: 1674.9 K, #E: 923.8 K Drug Development
-Hetionet (Jeong et al., 2022) 11 (e.g., biological process) 24 (e.g., disease–associates–gene) #N: 47.0 K, #E: 2250.2 K Drug Development
-LBD-COVID (Zhang et al., 2021b) 1 (i.e., concept) 1 (i.e., SemMedDB relation) #N: 131.4 K, #E: 1016.1 K Drug Development
-GP-KG (Gao et al., 2022) 7 (e.g., drug) 9 (e.g., disease–gene) #N: 61.1 K, #E: 1246.7 K Drug Development
-DRKF (Zhang and Che, 2021) 4 (e.g., drug) 43 (e.g., drug-disease) #N: 12.5 K, #E: 165.9 K Drug Development
-DDKG (Ghorbanali et al., 2023) 2 (i.e., drug & disease) 1 (e.g., drug-disease) #N: 551, #E: 2.7 K Drug Development
-Disease Ontology [^5] https://disease-ontology.org/ (Schriml et al., 2012) 1 (i.e., disease) 2 (e.g., subClassOf) #N: 11.2 K, #E: 8.8 K Clinical Decision Support
-DrugBank (Wishart et al., 2018) 4 (e.g., drug, pathway) 4 (e.g., drug-target) #N: 7.4 K, #E: 366.0 K Clinical Decision Support
-KnowLife (Ernst et al., 2014) 6 (e.g., genes) 14 (e.g., gene-diseases) #N: 2.9 M, #E: 11.4 M Clinical Decision Support
-PharmKG (Zheng et al., 2021) 3 (e.g., diseases) 3 (e.g., chemical-diseases) #N: 7601, #E: 500958 Clinical Decision Support
-ROBOKOP [^6] https://robokop.renci.org/ (Bizon et al., 2019) 54 (e.g., genes, drugs) 1064 (e.g., biolink, CHEBI) #N: 8.6M , #E: 130.4 M Clinical Decision Support
-iBKH [^7] https://github.com/wcm-wanglab/iBKH (Su et al., 2023) 11 (e.g., anatomy, disease) 18 (e.g., anatomy-gene) #N: 2.4 M, #E: 48.2 M Clinical Decision Support
-2.3. HKG Utilization
+HetioNet 81 11 (e.g., drug, disease) 24 (e.g., drug-disease) #N: 47.0 K, #E: 2.3 M Medicinal Chemistry
+DrKG 82 13 (e.g., disease, gene) 107 (e.g., disease-gene) #N: 97 K, #E: 5.8 M Medicinal Chemistry
+PrimeKG 11 10 (e.g., phenotypes) 30 (e.g., disease-phenotype) #N: 129.4 K, #E: 8.1 M Medicinal Chemistry
+Gene Ontology [^1] http://geneontology.org/ 35 3 (e.g., biological process) 4 (e.g., partOf) #N: 43 K, #E: 7544.6K Bioinformatics
+KEGG [^2] https://www.genome.jp/kegg/ 83 16 (e.g., pathway) 4 (e.g., partOf) #N: 48.5 M, #E: unknown Bioinformatics
+STRING [^3] https://string-db.org/ 84 1 (e.g., protein) 4 (e.g., interactions) #N: 67.6 M, #E: 20 B Bioinformatics
+Cell Ontology [^4] https://www.ebi.ac.uk/ols4/ontologies/cl 85 1 (i.e., cell type) 2 (e.g, subClassOf) #N: 2.7 K, #E: 15.9 K Bioinformatics
+GEFA 86 510 (e.g., kinases) 2 (e.g., drug-drug) #N: 0.5 K, #E: 30.1 K Drug Development
+Reaction 87 2 (e.g., reactant & normal) 19 (e.g., reaction paths) #N: 2192.7 K, #E: 932.2 K Drug Development
+ASICS 88 2 (e.g., reactant & product) 1 (e.g., reactions) #N: 1674.9 K, #E: 923.8 K Drug Development
+Hetionet 88 11 (e.g., biological process) 24 (e.g., disease–associates–gene) #N: 47.0 K, #E: 2250.2 K Drug Development
+LBD-COVID 89 1 (i.e., concept) 1 (i.e., SemMedDB relation) #N: 131.4 K, #E: 1016.1 K Drug Development
+GP-KG 90 7 (e.g., drug) 9 (e.g., disease–gene) #N: 61.1 K, #E: 1246.7 K Drug Development
+DRKF 91 4 (e.g., drug) 43 (e.g., drug-disease) #N: 12.5 K, #E: 165.9 K Drug Development
+DDKG 92 2 (i.e., drug & disease) 1 (e.g., drug-disease) #N: 551, #E: 2.7 K Drug Development
+Disease Ontology [^5] https://disease-ontology.org/ 36 1 (i.e., disease) 2 (e.g., subClassOf) #N: 11.2 K, #E: 8.8 K Clinical Decision Support
+DrugBank 93 4 (e.g., drug, pathway) 4 (e.g., drug-target) #N: 7.4 K, #E: 366.0 K Clinical Decision Support
+KnowLife 94 6 (e.g., genes) 14 (e.g., gene-diseases) #N: 2.9 M, #E: 11.4 M Clinical Decision Support
+PharmKG 95 3 (e.g., diseases) 3 (e.g., chemical-diseases) #N: 7601, #E: 500958 Clinical Decision Support
+ROBOKOP [^6] https://robokop.renci.org/ 96 54 (e.g., genes, drugs) 1064 (e.g., biolink, CHEBI) #N: 8.6M , #E: 130.4 M Clinical Decision Support
+iBKH [^7] https://github.com/wcm-wanglab/iBKH 48 11 (e.g., anatomy, disease) 18 (e.g., anatomy-gene) #N: 2.4 M, #E: 48.2 M Clinical Decision Support
+2.2 HKG Utilization
 Model-free Utilization.
-Various query languages can be used for KGs, such as SPARQL, Cypher, and GraphQL (Wang et al., 2020).
-These query languages allow users to query healthcare KGs using a standardized syntax, thus enabling users to retrieve, manipulate, and analyze data in a structured and consistent way.
-More complex applications can be further supported by graph queries.
-For instance, automatic healthcare question answering can be tackled by Natural Language Question-to-Query (NLQ2Query) approach (Kim et al., 2022), where natural language questions are first translated into executable graph queries and then answered by the query responses.
-HKGs can also be utilized as an up-to-date and trustworthy augmentation to large language models (LLMs) for many applications.
-Some pioneering studies (Liu et al., 2022c; Guu et al., 2020; Xu et al., 2023; Shi et al., 2023) show that retrieved knowledge triples can improve the reliability of LLMs in various knowledge-intensive tasks, by addressing the nonsensical or unfaithful generation.
-Moreover, KGs can be a useful tool for fact-checking (Tchechmedjiev et al., 2019; Vedula and Parthasarathy, 2021; Mayank et al., 2022) as they provide a structured representation of information that can be used to quickly and efficiently verify the accuracy of claims.
-Researchers have explored the utility of HKGs in identifying ingredient substitutions of food (Shirai et al., 2021), COVID-19 fact-checking (Mengoni and Yang, 2022), etc.
+Various query languages can be used for KGs, such as SPARQL, Cypher, and GraphQL 44.
+These query languages allow users to query HKGs using a standardized syntax, thus enabling users to retrieve, manipulate, and analyze data in a structured and consistent way.
+For instance, automatic healthcare question answering can be tackled by Natural Language Question-to-Query (NLQ2Query) approach 97, where natural language questions are first translated into executable graph queries and then answered by the query responses.
+HKGs can also be utilized as an up-to-date and trustworthy augmentation to LLMs for many applications.
+Some pioneering studies 98; 99; 100 show that retrieved knowledge triples can improve the reliability of LLMs in various knowledge-intensive tasks.
+Moreover, KGs can be a useful tool for fact-checking 101; 102; 103 as they provide a structured representation of information that can be used to quickly and efficiently verify the accuracy of claims.
+Researchers have explored the utility of HKGs in identifying ingredient substitutions of food 104, COVID-19 fact-checking 105, etc.
 Model-based Utilization.
 Utilizing HKGs in complex reasoning tasks often involves utilizing machine learning models.
-HKG embeddings (Yu et al., 2021; Su et al., 2022) have shown great potential to tackle these tasks.
-In particular, HKG embedding models are a class of machine learning models that aim to learn low-dimensional vector representations, or embeddings, of the entities and relations in a knowledge graph.
+HKG embeddings 106; 107 have shown great potential to tackle these tasks.
+In particular, HKG embedding models are a class of machine learning models that aim to learn low-dimensional vector representations of the entities and relations in a KG.
 After obtaining HGK embeddings, they can be plugged into any kind of deep neural network and further fine-tuned toward downstream objectives.
-On the other hand, symbolic logic models represent another prominent approach for KG reasoning due to their interpretability.
-More specifically, symbolic reasoning models first mine logical rules from existing knowledge by inductive logic programming (Muggleton, 1992), association rule mining (Galárraga et al., 2013), or Markov logic networks (Kok and Domingos, 2005).
+On the other hand, Symbolic logic models offer an interpretable approach to KG reasoning by mining logical rules from existing knowledge through techniques such as inductive logic programming 108, association rule mining 109, or Markov logic networks 110.
 These minded rules are used to infer new facts, make logical deductions and answer complex queries.
-Recently, researchers start to explore combining logical rules into KG embedding to further improve the generalization and performance of HKG reasoning (Alshahrani et al., 2017; Zhu et al., 2022a).
-3. Resources
-In this section, we compile a detailed resource overview of existing healthcare knowledge graphs to assist researchers and healthcare professionals in constructing and utilizing HKGs, organized in Table 1.
-We present key attribute information, including HKG name, node types, edge types, statistics, and their applications.
-Details and external links can be referred to at our public repository [^8] Resource: https://github.com/lujiaying/Awesome-HealthCare-KnowledgeBase.
+Recently, researchers start to explore combining logical rules into KG embedding to further improve the generalization and performance of HKG reasoning 111; 112.
+3 Applications
 […]
 L1/.style=draw=black, text width=2.5cm, L2/.style=draw=black, text width=4.2cm, L3/.style=draw=black, text width=4.5cm, L4/.style=draw=black, text width=4.5cm, {forest} for tree=draw,rounded corners,grow’=0,text width=2cm, text centered,edge+=-¿, forked edges, [Application, fill=blue!30, L1 [Basic Science Research, fill=blue!20, L2 [Medical Chemistry, fill=blue!10, L3] [Bioinformatics, fill=blue!10, L3] ] [Pharmaceutical Development, fill=blue!20, L2 [Drug Development, fill=blue!10, L3] [Clinical Trial, fill=blue!10, L3] ] [Clinical Decision Support, fill=blue!20, L2 [Intermediate Steps to Advance Prediction Models, fill=blue!10, L3] [Evidence Generation for Risk Prediction, fill=blue!10, L3] ] [Public Health, fill=blue!20, L2 [Epidemiology, fill=blue!10, L3] [Environmental Health, fill=blue!10, L3] [Policy and Management, fill=blue!10, L3] [Social and Behavioral, fill=blue!10, L3] ] ]
-Figure 4. Detailed taxonomy of the application section of healthcare knowledge graphs.
-4. Applications
-4.1. Basic Science Research
-Several previous biological terms can also be considered knowledge graphs, such as ontology (gene ontology, cell ontology, disease ontology), network (gene regulatory network), etc.
-We use the original biological terms as they are more popular according to historical reasons.
-4.1.1. Medicinal Chemistry
-Topics related to medicinal chemistry involve drug-drug interactions (DDIs) and drug-target interactions (DTIs), which will be discussed in this section.
-Drug-drug interactions (DDIs) refer to changes in the actions, or side effects, of drugs when they are taken at the same time or successively (Giacomini et al., 2007).
-In general, DDIs are a significant contributor to life-threatening adverse events (Su et al., 2022), and their identification is one of the key tasks in public health and drug development.
+Figure 5: Detailed taxonomy of the application section of healthcare knowledge graphs.
+3.1 Basic Science Research
+Biologists have built various ontologies (gene, cell, disease) and networks (gene regulatory, protein interaction) over the past decade, which are also KGs.
+They are commonly used in medicinal chemistry and bioinformatic research.
+3.1.1 Medicinal Chemistry
+Drug-drug interactions (DDIs) refer to changes in the actions, or side effects, of drugs when they are taken at the same time or successively 113.
+In general, DDIs are a significant contributor to life-threatening adverse events 107, and their identification is one of the key tasks in public health and drug development.
 The existence of diverse datasets on drug-drug interactions (DDIs) and biomedical KGs has enabled the development of machine-learning models that can accurately predict DDIs.
 Yu et al.
-(Yu et al., 2021) develop SumGNN, a model that includes a subgraph extraction module to efficiently extract relevant subgraphs from a KG, a self-attention-based summarization scheme to generate reasoning paths within the subgraph, and a multichannel module for integrating knowledge and data, resulting in significantly improved predictions of multi-typed DDIs.
+106 develop SumGNN, a model that efficiently extract relevant subgraphs from a KG to generate reasoning paths within the subgraph, resulting in significantly improved predictions of multi-typed DDIs.
 Su et al.
-(Su et al., 2022) propose DDKG, an attention-based KG representation learning framework that involves an encoder-decoder layer to learn the initial embeddings of drug nodes from their attributes in the KG.
+107 propose DDKG, an attention-based KG representation learning framework that involves an encoder-decoder layer to learn the initial embeddings of drug nodes from their attributes in the KG.
 Karim et al.
-(Karim et al., 2019) compare various techniques for generating KG embeddings with different settings and conclude that a combined convolutional neural network and LSTM network yields the highest accuracy when predicting drug-drug interactions (DDIs).
+114 compare various techniques for generating KG embeddings with different settings and conclude that a combined convolutional neural network and LSTM network yields the highest accuracy when predicting drug-drug interactions (DDIs).
 Dai et al.
-(Dai et al., 2021) propose a new KG embedding framework by introducing adversarial autoencoders based on Wasserstein distances and Gumbel-Softmax relaxation for DDI tasks.
+115 propose a new KG embedding framework by introducing adversarial autoencoders based on Wasserstein distances and Gumbel-Softmax relaxation for DDI tasks.
 Lin et al.
-(Lin et al., 2020) develop KGNN that resolves the DDI prediction by capturing drugs and their potential neighborhoods by mining associated relations in KG.
-Drug-target interactions (DTIs) is just as important as DDIs (Chen et al., 2016).
-Machine learning models can leverage knowledge graphs constructed from various types of interactions, such as drug-drug, drug-disease, protein-disease, and protein-protein interactions, to aid in predicting DTIs.
+116 develop KGNN that resolves the DDI prediction by capturing drugs and their potential neighborhoods by mining associated relations in KG.
+Drug-target interactions (DTIs) are as important as DDIs 117.
+Machine learning models can leverage KGs constructed from various types of interactions, such as drug-drug, drug-disease, protein-disease, and protein-protein interactions, to aid in predicting DTIs.
 For instance, Li et al.
-(Li et al., 2023) utilize the KG transfer probability matrix to redefine the drug-drug and target-target similarity matrix, thus constructing the final graph adjacent matrix to learn node representations by VGAE and augmenting them by utilizing dual Wasserstein Generative Adversarial Network with gradient penalty.
+118 utilize the KG transfer probability matrix to redefine the drug-drug and target-target similarity matrix, thus constructing the final graph adjacent matrix to learn node representations by utilizing dual Wasserstein Generative Adversarial Network.
 Zhang et al.
-(Zhang et al., 2021c) propose a new hybrid method for DTI prediction by first constructing DTI-related KGs and then employing graph representation learning model to obtain feature vectors of the KG.
+119 propose a new hybrid method for DTI prediction by first constructing DTI-related KGs and then employing graph representation learning model to obtain feature vectors of the KG.
 Wang et al.
-(Wang et al., 2022a) construct a knowledge graph of 29,607 positive drug-target pairs by DistMult embedding strategy, and propose a Conv-Conv module to extract features of drug-target pairs.
+120 construct a KG of 29,607 positive drug-target pairs by DistMult embedding strategy, and propose a Conv-Conv module to extract features of drug-target pairs.
 Ye et al.
-(Ye et al., 2021b) learn a low-dimensional representation for various entities in the KG, and then integrate the multimodal information via neural factorization machine.
-4.1.2. Bioinformatics Research
-In bioinformatics settings, a knowledge graph is a graph-based representation where nodes are biomedical entities (such as mutations, genes, proteins, metabolites, diseases, and biological pathways), and edges are their relationships (such as associations, interactions, and regulations).
-Through the integration, researchers can gain a more comprehensive understanding of complex biological processes and diseases.
-Multi-Omics Applications: In recent years, the field of multi-omics data analysis has become increasingly important for understanding complex biological systems.
-With the advancement of high-throughput technologies, researchers can generate large-scale and high-dimensional data from different omics fields, such as genomics, transcriptomics, proteomics, metabolomics, and epigenomics.
-More KG applications based on multi-omics data integration have emerged, aiming to provide new research methods to uncover the complex relationships between different omics layers and reveal biological systems’ underlying mechanisms.
-Knowledge graphs have been used to identify disease-associated mutations, genes, proteins, and metabolites by integrating multi-omics data with existing biological knowledge.
-This approach has led to the discovery of novel biomarkers and therapeutic targets for various diseases and the interpretation of the functional effects of genetic elements (Zhang et al., 2021a).
+121 learn a low-dimensional representation for various entities in the KG, and then integrate the multimodal information via neural factorization machine.
+3.1.2 Bioinformatics Research
+Multi-Omics Analysis has become increasingly important for understanding complex biological systems.
+With the advancement of high-throughput technologies, more KG applications based on multi-omics data integration have emerged, aiming to provide new research methods to uncover the complex relationships between different omics layers and reveal biological systems’ underlying mechanisms.
+KGs have been used to identify disease-associated mutations, genes, proteins, and metabolites by integrating multi-omics data with existing biological knowledge.
+This approach has led to the discovery of novel biomarkers and therapeutic targets for various diseases 122.
 Quan et al.
-built a comprehensive multi-relational knowledge graph called AIMedGraph, providing an interpretation of the impact of genetic variants on disease or treatment (Quan et al., [n.
-d.]).
+built a comprehensive multi-relational KG called AIMedGraph, providing an interpretation of the impact of genetic variants on disease or treatment 123.
 They curated detailed information about diseases, drugs, genetic variants, and the impact of genetic variations on disease development and drug treatment from multiple data resources.
-The entities integrated into AIMedGraph are connected by evidence-based relations and form a comprehensive gene–variant–disease–drug–trial–reference knowledge network.
-AIMedGraph uses the Adamic-Adar algorithm to predict new relations between entities based on shared neighbors and their proximity in the knowledge network.
-GenomicsKG is a knowledge graph to analyze and visualize multi-omics data.
+GenomicsKG is a KG to analyze and visualize multi-omics data.
 GenomicsKG can be used to improve drug development based on clinical genomics correlations and personalized drug customization in the extended version based on interactive relationships.
-It also provides multi-dimensional visualization, linked functional knowledge graphs, and reporting for clinical genomics.
-Single-Cell Analysis: Cells are fundamental and essential units of living organisms.
-With high-throughput sequencing technologies advancing to measure genomic profiles in a single-cell resolution, cell functions (inside cells) and cell-cell interactions (between cells) are revealed (Linnarsson and Teichmann, 2016).
-When diving into the functions of cells, gene regulatory mechanisms can be a critical factor.
-Gene regulatory mechanisms control the expression of genes, which can affect cell differentiation, response to stimuli, disease progression, etc.
-It can be visualized as gene regulatory networks (GRNs) which depict the interactions between genes and their regulators.
-Traditional approaches to constructing the GRN are based on gene knockdown or knockout experiments, which can be time-consuming, labor-intensive, and costly.
-In addition, these experiments are limited to only one or a few genes which neglects gene-gene interactions.
-In contrast, single-cell sequencing data could provide whole-genome scale measurements in each individual cell with comprehensive information such as gene expression, transcription factors (TF) binding sites, DNA methylation, epigenetic modifications, etc.
-By mining these publicly available data, it is possible to reveal the underlying and universal GRN biomedical researchers to understand biological processes better.
-For example, GRNdb provides detailed regulon and TF-target pairs information from different human and mouse tissues under different conditions by analyzing existing sequencing data (Fang et al., 2021).
-GenomicKB integrates existing datasets and genome annotations and formulates the data into a knowledge graph to emphasize the relationships among genomic entries (Feng et al., 2023).
-We can foresee that with more data being generated and collected, one generalized GRN (gene regulatory knowledge graph) along with cell-type-specific GRNs can be reconstructed and used to help reveal the underlying mechanisms of gene expression, biological processes, and disease progression.
-Besides cell-type-specific GRN, single-cell sequencing data can also be used to infer cell-cell interactions, which play a critical role in understanding cell cycles, cell fate decisions, tissue development, immune response, etc.
-Among several approaches through which cells can interact, cell-cell communication or cell signaling is of the most interest (Eltzschig et al., 2006) as the physical distance between cells does not limit it.
-Cells can communicate with each other by sending signaling molecules called ligands and receiving through receptors located on cell surfaces.
-Several methods have been developed to implicitly measure cell-cell communication by modeling and scoring ligand-receptor interactions (LRIs) from single-cell sequencing data (Efremova et al., 2020; Jin et al., 2021).
-However, these methods do not utilize the spatial information of cells.
-Recent advancements in spatial sequencing techniques provide not only genomic measurements but also complementary information on the spatial coordinates of each cell, where the spatial coordinates indicate the probabilities of cells communicating with each other as cells with closer distances tend to interact more.
-A recent benchmark study on evaluating cell-cell communication inference methods based on spatial distances (Liu et al., 2022a) while another benchmark study also considers other biological factors such as cytokine activities and receptor protein abundance (Dimitrov et al., 2022).
-With the availability of spatial transcriptomics data, many methods have been developed to leverage colocalization information to infer LRIs.
-Although many LRIs databases (Efremova et al., 2020; Shao et al., 2021) have been constructed and applied to infer cell-cell communication, only recently, SpaTalk (Shao et al., 2022) integrates CellTalkDB, Kyoto Encyclopedia of Genes and Genomes (KEGG) pathways, Reactome and TFs from AnimalTFDB to construct a ligand-receptor-target knowledge graph to help improve the inference of cell-cell communication.
-The success of SpaTalk provides an unprecedented view of integrating existing networks into a knowledge graph and shows that correctly utilizing the knowledge graph could provide helpful information in cell-cell communication tasks.
-4.2. Pharmaceutical Research Development
-4.2.1. Drug Development
+It also provides multi-dimensional visualization, linked functional KGs, and reporting for clinical genomics.
+Single-Cell Analysis focuses on cells as the fundamental and essential units of living organisms.
+With high-throughput sequencing technologies advancing to measure genomic profiles in a single-cell resolution, cell functions (inside cells) and cell-cell interactions (between cells) are revealed 124.
+Gene regulatory mechanisms, which control gene expression and affect processes like cell differentiation and disease progression, are key to understanding these functions.
+Traditional gene knockdown experiments are time-consuming and limited, but single-cell sequencing provides genome-wide data, including gene expression, transcription factor binding, DNA methylation, and epigenetic modifications.
+This allows researchers to uncover gene regulatory networks (GRNs) that enhance our understanding of biological processes.
+Databases like GRNdb and GenomicKB integrate sequencing data and annotations to provide insights into gene regulation across tissues 125; 126, which will improve as more data becomes available.
+3.2 Pharmaceutical Research Development
+3.2.1 Drug Development
 Drug development is the process of identifying novel chemical compounds that can effectively treat or alleviate human diseases.
 Before the drug can be designated as a final product for clinical use, several critical steps need to be undertaken from the initial target identification, chemical synthesis and clinical trials.
-The whole process typically spans over a decade and involves expenses of approximately one billion dollars (Adams and Brantner, 2006), yet it is characterized by a low success rate for clinical approval (DiMasi et al., 2015).
-With advancements in chemistry and the availability of vast chemical libraries, it is possible to reduce the costs of drug development by extracting and integrating valuable insights from existing data and using computer algorithms to accelerate the drug design process (Du et al., 2021; Zeng et al., 2022; Wang et al., 2022c; Pan et al., 2022b; Wang et al., 2022b; Nicolaou et al., 2007; Nicolaou and Pattichis, 2006).
-Despite the growing trend of computer-assisted drug discovery, a key question remains regarding how to effectively integrate data and extract valuable insights from the vast chemical dataset.
-To approach this question, knowledge graphs (KGs) have been employed for drug discovery due to their various advantages (Zeng et al., 2022): (1) In contrast to traditional methods that capture only one type of relationship, KGs are capable of providing heterogeneous information that includes diverse entities (e.g., scaffolds, proteins, and genes); (2) In addition, KGs are capable of handling multiple types of relationships between various types of entities, such as drug-target pairs; and (3) KGs can provide unstructured semantic relationships between entities.
-In such graphs, entities are represented as nodes while their relationships are represented as edges, by which complex relations in biochemical systems can be easily handled.
-In general, the field of drug development encompasses two main areas: drug design and drug repurposing (also known as drug repositioning).
-Drug design is the process of creating novel and diverse drug molecules with desirable pharmaceutical properties (Jing et al., 2018; Fu et al., 2021; Fu and Sun, 2022), whereas drug repurposing identifies new uses for existing approved drugs that were originally developed for a different indication (Pan et al., 2022a; Huang et al., 2020).
-Drug Design.
-KGs are widely employed in drug design, particularly in generating novel molecules that are promising drug candidates for various diseases (Ranjan et al., 2022; Li and Chen, 2022).
+The whole process typically spans over a decade and involves expenses of approximately one billion dollars 127, yet it is characterized by a low success rate for clinical approval 128.
+Drug Design is an area where KGs are commonly used, particularly for generating novel molecules that are promising drug candidates for various diseases 86; 87.
 Ranjan et al.
-(Ranjan et al., 2022) utilize Gated Graph Neural Network (GGNN) to generate novel molecules that target the coronavirus (i.e., SARS-CoV-2) (Hasöksüz et al., 2020) and integrate KGs into their approach to reduce the search space.
+86 utilize Gated Graph Neural Network (GGNN) to generate novel molecules that target the coronavirus (i.e., SARS-CoV-2) 129 and integrate KGs into their approach to reduce the search space.
 Specifically, KGs were leveraged to discard non-binding molecules before inputting them into the Early Fusion model, thus optimizing the efficiency of the drug design process.
@@
 For instance, Li et al.
-(Li and Chen, 2022) utilize reaction KGs to construct classification models for compound synthetic accessibility.
+87 utilize reaction KGs to construct classification models for compound synthetic accessibility.
 By leveraging KGs that capture information about reactions, including reaction types, substrates, and reaction conditions, they can train machine learning models that could predict the synthetic accessibility of compounds.
 Jeong et al.
-(Jeong et al., 2022) introduce an intelligent system that integrates generative exploration and exploitation of reaction knowledge base to support synthetic path design.
-Drug Repurposing.
-Compared to drug design, KGs are more commonly utilized to expedite the drug re-purposing process (Zhu et al., 2020; MacLean, 2021; Himmelstein et al., 2017; Zhang et al., 2021b; Gao et al., 2022; Xu et al., 2019; Zhang and Che, 2021; Ghorbanali et al., 2023).
-Many applications on drug re-purposing that utilize KGs are primarily focused on link prediction tasks (MacLean, 2021).
-To re-purpose promising drug candidates for new indications, many methods employ predictive models that focus on predicting drug-treats-disease relationships within pharmacological knowledge graphs KGs.
-Himmelstein et al.
-(Himmelstein et al., 2017) use a degree-normalized pathway model on the hetionet KG, which includes genes, diseases, tissues, pathophysiologies, and multimodal edges, to identify potentially repurposable drugs for epilepsy.
+88 introduce an intelligent system that integrates generative exploration and exploitation of reaction knowledge base to support synthetic path design.
+Drug Repurposing has often been expedited by the utilization of KGs 130; 131; 47; 89; 90; 132; 91; 92.
+Many applications on drug re-purposing that utilize KGs are primarily focused on link prediction tasks 131.
+To re-purpose promising drug candidates for new indications, many methods employ predictive models that focus on predicting drug-treats-disease relationships within pharmacological KGs.
 Xu et al.
-(Xu et al., 2019) develop a multi-path random walk model on a network that incorporates gene-phenotype associations, protein-protein interactions, and phenotypic similarities for training and prediction purposes.
+132 develop a multi-path random walk model on a network that incorporates gene-phenotype associations, protein-protein interactions, and phenotypic similarities for training and prediction purposes.
 Zhang et al.
-(Zhang et al., 2021b) introduce an integrative and literature-based discovery model for identifying potential drug candidates from COVID-19-focused research literature, including PubMed and other relevant sources.
+89 introduce an integrative and literature-based discovery model for identifying potential drug candidates from COVID-19-focused research literature, including PubMed and other relevant sources.
 Gao et al.
-(Gao et al., 2022) construct a knowledge graph (KG) by integrating multiple genotypic and phenotypic databases.
+90 construct a KG by integrating multiple genotypic and phenotypic databases.
 They then learn low-dimensional representations of the KG and utilize these representations to infer new drug-disease interactions, providing insights into potential drug repurposing opportunities.
-Zhang and Che (Zhang and Che, 2021) introduce a model for drug re-purposing in Parkinson’s disease that leverages a local medical knowledge base incorporating accurate knowledge along with medical literature containing novel information.
+Zhang and Che 91 introduce a model for drug re-purposing in Parkinson’s disease that leverages a local medical knowledge base incorporating accurate knowledge along with medical literature containing novel information.
 Ghorbanali et al.
-(Ghorbanali et al., 2023) present the DrugRep-KG method, which utilizes a KG embedding approach for representing drugs and diseases in the process of drug repurposing.
-4.2.2. Clinical Trial
+92 present the DrugRep-KG method, which utilizes a KG embedding approach for representing drugs and diseases in the process of drug repurposing.
+3.2.2 Clinical Trial
 The major goal of clinical trials is to assess the safety and effectiveness of drug molecules on human bodies.
 A novel drug molecule needs to pass three phases of clinical trials before it is approved by the Food and Drug Administration (FDA) and enters the drug market.
-The whole process is prohibitively time-consuming and expensive, costing 7-11 years and two billion dollars on average (Martin et al., 2017).
-Clinical Trial Optimization targets identifying eligible patients for clinical trials based on their medical history and health conditions (Rivera et al., 2020; He et al., 2020).
-Recently, with massive electronic health records (EHR) data and trial eligibility criteria (EC), data-driven methods have been studied to automatically assign appropriate patients for clinical trials (Yuan et al., 2019; Tseo et al., 2020; Liu et al., 2021).
+The whole process is prohibitively time-consuming and expensive, costing 7-11 years and two billion dollars on average 133.
+Clinical Trial Optimization targets identifying eligible patients for clinical trials based on their medical history and health conditions 134; 135.
+Recently, with massive electronic health records (EHR) data and trial eligibility criteria (EC), data-driven methods have been studied to automatically assign appropriate patients for clinical trials 136; 137; 138.
 However, it is often hard to fully capture and represent the complex knowledge present in unstructured ECs and EHR data, as ECs may only provide general disease concepts.
@@
 To better capture the interactions among different medical concepts from EHR records and ECs, Gao et al.
-(Gao et al., 2020) enhance patient records with hierarchical taxonomies to align medical concepts of varying granularity between EHR codes and ECs.
+139 enhance patient records with hierarchical taxonomies to align medical concepts of varying granularity between EHR codes and ECs.
 Besides, Fu et al.
-(Fu et al., 2022) leverage additional knowledge-embedding modules along with drug pharmacokinetic and historical trial data to improve the patient trial optimization process, and Wang et al.
-(Wang et al., 2023) leverage the knowledge graphs to learn static trial embedding and further designed meta-learning module to generalize well over the imbalanced clinical trial distribution.
-4.3. Clinical Decision Support
-Nowadays, abundant Electronic Health Record (EHR) data enables better computational models for accurate diagnoses and treatments.
-EHR contains essential patient information such as disease diagnoses, prescribed medications, and test results.
-Due to this valuable information, EHRs are extensively utilized to identify patterns in patient health and assist healthcare providers in making informed clinical decisions.
-However, the sparsity of EHR data typically allows for only a small fraction of medical codes to be learned effectively, thereby restricting the ability of deep learning approaches.
-To overcome this drawback, knowledge graphs have been applied to incorporate prior medical knowledge for these deep learning models, which augment the representation of medical codes to better support the downstream prediction tasks.
-4.3.1. Intermediate Steps to Advance Prediction Models
-ICD Coding aims to extract diagnosis and procedure codes from clinical notes, which often consisted of raw text (Mullenbach et al., 2018; Zhang et al., 2020; Vu et al., 2021; Dong et al., 2022).
-It is often challenging, as the size of the candidate target codes can be large and the distribution of the codes is often long-tailed (Kim and Ganapathi, 2021).
+140 leverage additional knowledge-embedding modules along with drug pharmacokinetic and historical trial data to improve the patient trial optimization process, and Wang et al.
+141 leverage the KGs to learn static trial embedding and further designed meta-learning module to generalize well over the imbalanced clinical trial distribution.
+3.3 Clinical Decision Support
+Electronic Health Record (EHR) contains essential patient information such as disease diagnoses, prescribed medications, and test results.
+However, the sparsity of EHR data largely restricts the ability of deep learning approaches.
+To overcome this drawback, KGs have been applied to incorporate prior medical knowledge for these deep learning models to better support the downstream prediction tasks.
+3.3.1 Intermediate Steps to Advance Prediction Models
+ICD Coding aims to extract diagnosis and procedure codes from clinical notes which are often raw texts 142; 143; 144; 145.
+It is often challenging, as the size of the candidate target codes can be large and the distribution of the codes is often long-tailed 146.
 To overcome this, Xie et al.
-(Xie et al., 2019) and Cao et al.
-(Cao et al., 2020a) propose to leverage knowledge graphs as distant supervision (Min et al., 2013; Liang et al., 2020), and inject the label information via structured knowledge graph propagation by leveraging graph convolution networks (Kipf and Welling, 2016) to learn the correlations among medical codes.
+147 and Cao et al.
+148 propose to leverage KGs as distant supervision 149; 58, and inject the label information via structured KG propagation by leveraging graph convolution networks 150 to learn the correlations among medical codes.
 Besides, Lu et al.
-(Lu et al., 2020) propose to leverage knowledge graphs and the co-occurrence graph among clinical nodes simultaneously with a knowledge aggregation module to boost the ICD coding performance.
+151 propose to leverage KGs and the co-occurrence graph among clinical nodes simultaneously with a knowledge aggregation module to boost the ICD coding performance.
 Ren et al.
-(Ren et al., 2022) design a learning curriculum based on the hierarchical structure of the code to address the highly imbalanced label distribution issue and balance between frequent and rare labels.
+152 design a learning curriculum based on the hierarchical structure of the code to address the highly imbalanced label distribution issue and balance between frequent and rare labels.
 Overall, injecting additional knowledge with graph neural networks offers a way to mitigate the imbalanced label distribution issue and thus better.
-Entity and Relation Extraction from Health Records.
-Health records contain rich unstructured or semi-structured data, making it difficult for clinicians to access and analyze relevant information.
-Entity and relation extraction helps convert this unstructured text into structured data that can be more easily processed, understood, and utilized.
+Entity and Relation Extraction helps convert the rich unstructured or semi-structured data in health records text into structured data that can be more easily processed, understood, and utilized by clinicians and algorithms.
 Specifically, entity extraction aims to identify entity mentions from clinical-free texts.
 There are two key steps for entity extraction, i.e., named entity recognition (NER) and disambiguation (NED).
-By leveraging additional knowledge graphs, Varma et al.
-(Varma et al., 2021) transfer structural knowledge from the knowledge base to the medical domain, improving rare entities’ disambiguation accuracy.
-Yuan et al.
-(Yuan et al., 2023) inject additional knowledge from the knowledge graphs for entity linking and proposed two additional strategies, namely Post-pruning and Thresholding, to improve the efficiency and remove the effect of unlinkable entity mention.
+By leveraging additional KGs, Yuan et al.
+153 inject additional knowledge from the KGs for entity linking and proposed post-pruning and Thresholding to improve the efficiency and remove the effect of unlinkable entity mention.
 Fries et al.
-(Fries et al., 2021) leverage clinical ontologies to provide weak supervision sources to create additional training data for clinical entity disambiguation.
+73 leverage clinical ontologies to provide weak supervision sources to create additional training data for clinical entity disambiguation.
 Besides, relation extraction aims to identify and classify relationships between entities in unstructured text, which facilitates understanding complex biological processes, drug interactions, and disease mechanisms.
-To incorporate the external knowledge graph, several works (Fei et al., 2021; Roy and Pan, 2021) proposed additional post-training steps to align the language models with biomedical knowledge.
+To incorporate the external KG, several works 154; 155 proposed additional post-training steps to align the language models with biomedical knowledge.
 Hong et al.
-(Hong et al., 2021) construct embeddings for a wide range of codified concepts from EHRs to identify relevant features related to a disease of interest, and Lin et al.
-(Lin et al., 2022) design a co-training scheme to jointly learn from text and knowledge graphs for extracting and classifying disease-disease relations.
-In summary, fusing knowledge graphs with language models can flexibly accommodate missing data types and bring additional performance gains, especially for those rare entities and relations.
-4.3.2. Evidence Generation for Risk Prediction Models
+156 construct embeddings for a wide range of codified concepts from EHRs to identify relevant features related to a disease of interest, and Lin et al.
+157 design a co-training scheme to jointly learn from text and KGs for extracting and classifying disease-disease relations.
+In summary, fusing KGs with language models can flexibly accommodate missing data types and bring additional performance gains, especially for those rare entities and relations.
+3.3.2 Evidence Generation for Risk Prediction Models
 Disease Prediction aims to predict the potential diseases of a given patient with his past clinical records.
-To assist the diagnosis with additional knowledge, GRAM (Choi et al., 2017) and KAME (Ma et al., 2018) utilize a medical ontology (Dubberke et al., 2006) where the leaf nodes are the medical codes found in EHR data, and their ancestors are more general categories.
+To assist the diagnosis with additional knowledge, GRAM 158 and KAME 159 utilize a medical ontology 160 where the leaf nodes are the medical codes found in EHR data, and their ancestors are more general categories.
 By incorporating information from medical ontologies into deep learning models via neural attention, these approaches learn better embeddings for different medical concepts to alleviate the data scarcity bottleneck.
-(Yin et al., 2019; Zhang et al., 2019) further consider the domain-specific knowledge graph KnowLife (Ernst et al., 2015) to enrich the embeddings of medical entities with their neighbors on the knowledge graph.
-These approaches mainly directly update the embeddings of different concepts to improve the feature learning, but may be at the risk of ignoring the high-level order information from the knowledge graph.
+161; 162 further consider the domain-specific KG KnowLife 163 to enrich the embeddings of medical entities with their neighbors on the KG.
+These approaches mainly directly update the embeddings of different concepts to improve the feature learning, but may be at the risk of ignoring the high-level order information from the KG.
 To tackle this drawback, Ye et al.
-(Ye et al., 2021a) explicitly exploit paths in KG from the observed symptoms to the target disease to model the personalized information for diverse patients with a relational-guided attention mechanism.
+164 explicitly exploit paths in KG from the observed symptoms to the target disease to model the personalized information for diverse patients.
 Xu et al.
-(Xu et al., 2021) design a self-supervised learning approach to pre-train a graph attention network for learning the embedding of medical concepts and completing the knowledge graph simultaneously.
+165 design a self-supervised learning approach to pre-train a graph attention network for learning the embedding of medical concepts and completing the KG simultaneously.
 These approaches better harness the structure information, and often lead to better performance than the pure embedding-based knowledge integration techniques.
-Treatment Recommendation aims to recommend personalized medications to patients based on their individual health conditions, which can help physicians select the most effective medications for their patients, and improve treatment outcomes (Zhang et al., 2017; Bhoi et al., 2021; Shang et al., 2019b).
+Treatment Recommendation aims to recommend personalized medications to patients based on their individual health conditions, which can help physicians select the most effective medications for their patients, and improve treatment outcomes 166; 167; 168.
 To effectively exploit external knowledge, Shang et al.
-(Shang et al., 2019a) use drug ontologies to design additional pretraining loss and directly improve the representation of drugs, and several studies (Wu et al., 2022a; Tan et al., 2022) attempt to extract the additional drug interaction graphs to model the negative side effects of specific drug pairs and reduce the possibility of recommending negative drug-drug interaction combinations.
+169 use drug ontologies to design additional pretraining loss and directly improve the representation of drugs, and several studies 170; 171 attempt to extract the additional drug interaction graphs to model the negative side effects of specific drug pairs and reduce the possibility of recommending negative drug-drug interaction combinations.
 Besides, Wu et al.
-(Wu et al., 2022b) leveraged ontologies to improve the drug representations, and facilitates drug recommendation under a more challenging few-shot setting.
-4.4. Public Health
+172 leveraged ontologies to improve the drug representations, and facilitates drug recommendation under a more challenging few-shot setting.
+3.4 Public Health
 Public Health research can significantly benefit from HKGs.
-Knowledge graphs can help organize, structure, and formalize extensive information from diverse and heterogeneous sources.
+KGs can help organize, structure, and formalize extensive information from diverse and heterogeneous sources.
 This allows researchers to analyze data, reason about factors, and make decisions on a larger scale.
 Epidemiology.
-The field of epidemiology has seen an increase in the use of knowledge graphs to analyze and understand the spread of diseases.
-A study conducted by Gao et al.
-(Gao et al., 2023) analyzes the research hotspots and development trends of wastewater-based epidemiology (WBE) using knowledge graphs constructed from nearly 900 papers.
+The field of epidemiology has seen an increased use of KGs to analyze and understand the spread of diseases.
+A study by Gao et al.
+173 analyzes the research and development trends of wastewater-based epidemiology (WBE) using KGs constructed from nearly 900 papers.
 Domingo-Fernández et al.
-(Domingo-Fernández et al., 2021) create the COVID-19 Knowledge Graph, a comprehensive cause-and-effect network constructed from the scientific literature on the coronavirus.
+174 create the COVID-19 KG, a comprehensive cause-and-effect network constructed from the scientific literature on the coronavirus.
 Additionally, Turki et al.
-(Turki et al., 2022) use knowledge graphs to assess and validate the portion of Wikidata related to COVID-19 epidemiology using an automatable task set.
+175 use KGs to assess and validate the portion of Wikidata related to COVID-19 epidemiology using an automatable task set.
 Pressat Laffouilhère et al.
-(Pressat Laffouilhère et al., 2022) develop OntoBioStat, a domain ontology related to covariate selection and bias in biostatistics, which can help interpret significant statistical associations between variables.
+176 develop OntoBioStat, a domain ontology related to covariate selection and bias in biostatistics, which can help interpret significant statistical associations between variables.
 Environmental Health.
 Fecho et al.
-(Fecho et al., 2021) develop ROBOKOP, a biomedical knowledge graph-based system, to validate associations between workplace chemical exposures and immune-mediated diseases.
+177 develop ROBOKOP, a biomedical KG-based system, to validate associations between workplace chemical exposures and immune-mediated diseases.
 Wolffe et al.
-(Wolffe et al., 2020) propose using knowledge graphs in systematic evidence mapping in environmental health.
+178 propose using KGs in systematic evidence mapping in environmental health.
 This approach overcomes the limitations of rigid data tables by offering a more suitable model for handling the highly connected and complex nature of environmental health data.
@@
 Wu et al.
-(Wu, 2021) have analyzed the COVID-19 epidemic situation using a knowledge graph of patient activity.
+179 have analyzed the COVID-19 epidemic situation using a KG of patient activity.
 This method enables in-depth study of the transmission process, analysis of key nodes, and tracing of activity tracks.
 Meanwhile, Yu et al.
-(Yu et al., 2022) develop a chronic management system, which combines knowledge graphs and big data to optimize the management of chronic diseases in children.
+180 develop a chronic management system, which combines KGs and big data to optimize the management of chronic diseases in children.
 This system enhances treatment and resource utilization while conforming to the requirements of the Chronic Care Model.
 Social and Behavioral Health.
-There have been several interesting research studies.
 Cao et al.
-(Cao et al., 2020b) build a high-level suicide-oriented knowledge graph combined with deep neural networks for detecting suicidal ideation on social media platforms.
+181 build a high-level suicide-oriented KG combined with deep neural networks for detecting suicidal ideation on social media platforms.
 Also, Liu et al.
-(Liu et al., 2020) conduct a bibliometric analysis of driver behavior research.
+182 conduct a bibliometric analysis of driver behavior research.
 Additionally, Wang et al.
-(Wang et al., 2022d) create an analysis framework for interpreting causal associations in emotional logic.
-They introduce a knowledge graph into appraisal theories, improving human emotional inference.
-5. Promise and Outlook
-The potential impact of comprehensive and fine-grained HKGs on biomedical research and clinical practice is significant.
+183 create an analysis framework for interpreting causal associations in emotional logic.
+They introduce a KG into appraisal theories, improving human emotional inference.
+Table 2: Main challenges in different types of problems.
+Types of Problems Main Challenges
+Basic Science Research - Limited integration of KGs with LLMs for structural and functional predictions of molecules and genes 184.
+- Difficulty in representing high-dimensional data (e.g., protein structures) in KGs 185.
+- Lack of domain-specific fine-tuning of LLMs for bioinformatics tasks 186.
+Pharmaceutical Development - Incomplete or inconsistent drug-related knowledge representation in KGs 187.
+- Challenges in combining structured KG data with unstructured LLM outputs (e.g., literature mining) for assisting drug discovery 92.
+Clinical Decision Support - Difficulty in ensuring explainability and reliability of predictions 188.
+- Lack of harmonization between real-time clinical data and static KGs for dynamic decision-making 32.
+Public Health - Limited ability to model causal relationships and complex interactions in public health domains using KGs 189.
+- Difficulty in integrating population-level data with KGs for real-time surveillance. 190
+4 Challenge, Promise, and Outlook
+Current Challenges for Adopting KGs to Biomedical Research.
+Injecting KGs into clinical problems faces significant challenges across domains.
+In basic science research, the integration of KGs and LLMs for molecular predictions is limited by difficulties in representing high-dimensional data like protein structures and a lack of domain-specific fine-tuning 184; 185; 186.
+In pharmaceutical development, inconsistent drug-related knowledge representation and challenges in combining KGs with unstructured LLM outputs hinder drug discovery 187; 92.
+Clinical decision support struggles with ensuring explainability, reliability, and harmonizing real-time clinical data with static KGs for adaptive decision-making 188; 32.
+Public health faces limitations in modeling causal relationships and integrating population-level data with KGs for real-time surveillance due to scalability and standardization issues 189; 190.
+Addressing these barriers is vital for leveraging KGs to advance clinical and public health outcomes.
+Future Directions The potential impact of comprehensive and fine-grained HKGs on biomedical research and clinical practice is significant.
 By integrating vast amounts of biomedical knowledge from multiple domains, HKGs can facilitate the discovery of new disease mechanisms and the identification of novel drug targets.
 They also help to enable personalized medicine by identifying patient subgroups with shared disease mechanisms.
-The recent success of large language models (a.k.a.
-foundation models) such as ChatGPT offers promising opportunities in capturing such semantics from the biomedical context (Agrawal et al., 2022; Singhal et al., 2022; Nath et al., 2022; Moor et al., 2023), enabling the construction of unprecedentedly comprehensive HKGs.
+The recent success of LLMs such as ChatGPT offers promising opportunities in capturing such semantics from the biomedical context 72; 191; 192; 193, enabling the construction of unprecedentedly comprehensive HKGs.
 In turn, HKGs also help improve LLMs by providing accurate and contextualized knowledge to regularize the generated content.
 This is particularly useful in evaluating LLMs in biomedical applications and addressing the problem of hallucination in critical areas.
-6. Conclusion
-Healthcare knowledge graphs have emerged as a promising approach for capturing and organizing medical knowledge in a structured and interpretable way.
-This comprehensive review paper provides an overview of the current state of HKGs, including their construction, utilization models, and applications in healthcare.
-Furthermore, the paper discusses the potential future developments of HKGs.
+For developers of novel biomedical informatics methods, this symbiotic relationship between HKGs and LLMs presents both opportunities and challenges.
+One key area for innovation is the design of frameworks that allow real-time updating of HKGs with the latest clinical and research data, enabling dynamic decision-making.
+Developers must also address the computational and representational challenges of integrating high-dimensional data, such as genomic sequences and protein structures, into HKGs while ensuring semantic consistency.
+Moreover, constructing HKGs that incorporate causal reasoning capabilities could significantly enhance their utility in understanding complex disease mechanisms and predicting therapeutic outcomes.
+Another promising avenue lies in leveraging fine-tuned LLMs to extract domain-specific insights from unstructured biomedical literature and seamlessly integrate them into HKGs.
+This process, however, requires rigorous validation pipelines to ensure the accuracy and relevance of the extracted information.
+Developers should also explore strategies for aligning LLM outputs with ontologies and existing knowledge schemas to maximize interoperability and reuse.
+Finally, as HKGs are increasingly utilized to support clinical decision-making, developers must prioritize explainability and fairness in their design.
+This includes developing visualization tools and interactive interfaces that allow clinicians and researchers to interpret the outputs of HKG-powered systems effectively.
+Similarly, integrating multilingual capabilities and context-aware reasoning into HKGs will be essential for addressing global healthcare challenges and ensuring equitable access to cutting-edge biomedical insights.
+5 Conclusion
+HKGs have emerged as a promising approach for capturing and organizing medical knowledge in a structured and interpretable way.
+This comprehensive review paper provides an overview of the current state of HKGs, including their construction, modeling, and applications in healthcare.
+Furthermore, the paper discusses potential future developments of HKGs.
 In conclusion, HKGs have played a significant role in advancing health research.
@@
 We hope that our comprehensive review of this field offers a helpful perspective for future reference.
+Acknowledgements
+This research was partially supported by the National Science Foundation under Award Number 2319449 and Award Number 2312502, as well as the National Institute Of Diabetes And Digestive And Kidney Diseases of the National Institutes of Health under Award Number K25DK135913.
+Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of the National Science Foundation, National Institutes of Health, or the U.S.
+government.
+The authors wish to thank the editors and reviewers for their valuable efforts and suggestions.